### 28 Nov 2023 v1

# Cole: A Hierarchical Generation Framework for Graphic Design

Abstract
Graphic design, which has been evolving since the 15th century, plays a crucial role in advertising. The creation of high-quality designs demands creativity, innovation, and lateral thinking. This intricate task involves understanding the objective, crafting visual elements such as the background, decoration, font, color, and shape, formulating diverse professional layouts, and adhering to fundamental visual design principles. In this paper, we introduce Cole 111In 1843, Henry Cole introduced the world’s first commercial Christmas card [37]. We commemorate his contributions with the Cole system., a hierarchical generation framework purposed to comprehensively address these challenges. This Cole system can transform a straightforward intention prompt into a high-quality graphic design, while also supporting flexible editing based on user input. Examples of such input might include directives like “design a poster for Hisaishi’s concert.” The key insight is to dissect the complex task of text-to-design generation into a hierarchy of simpler sub-tasks, each addressed by specialized models working collaboratively. The results from these models are then consolidated to produce a cohesive final output. Our hierarchical task decomposition can streamline the complex process and significantly enhance generation reliability. Our Cole system consists of multiple fine-tuned Large Language Models (LLMs), Large Multimodal Models (LMMs), and Diffusion Models (DMs), each specifically tailored for a design-aware text or image generation task. Furthermore, we construct the DesignerIntention benchmark to highlight the superiority of our Cole over existing methods in generating high-quality graphic designs from user intent. We perceive our Cole as an important step towards addressing more complex visual design generation tasks in the future.
1 Introduction
The recent advancement in the quality of natural image generation has been remarkable, elevating it to the level of professional photography. This leap forward can be attributed to the development of technologies such as Imagen [7], SDXL [26], and DALLE3 [2]. The pivotal factors contributing to these advancements include the use of powerful Large Language Model (LLM) as a text encoder, the scaling up of training datasets, increased model complexity, the design of advanced sampling strategies, and the enhancement of data quality, among others. We believe that now is the opportune moment to redirect our efforts towards more professional image generation, particularly in the area of graphic design generation considering its pivotal roles in advertising, branding, and marketing.
Graphic design [36], a professional discipline, leverages the power of visual communication to convey targeted messages to specific social groups, with clear objectives. It is a realm that necessitates creativity, innovation, and lateral thinking. Typically, graphic design employs either manual or digital tools to amalgamate text and graphics, thus creating compelling visual narratives. Its primary goal is to structure information, give substance to ideas, and infuse expression and emotion into artifacts that chronicle human experiences. Graphic design often utilizes the artistry of typography, text composition, ornamentation, and imagery to convey thoughts, emotions, and attitudes that transcend the capabilities of language alone. In essence, it demands a high degree of creativity, innovation, and lateral thinking to produce top-tier designs.
The most recent study [18] has empirically shown that the groundbreaking DALLE3 exhibits impressive capabilities in generating high-quality design images, marked by visually striking graphics and layouts, as illustrated in Figure 2. However, these images are not without their flaws. They continue to confront several critical challenges, such as mis-rendered visual text, often resulting in missing or extra characters, a phenomenon also highlighted in [2]. Furthermore, editing these generated images demands complex operations such as segmentation, erasing, inpainting, and others, owing to their inherently uneditable nature. Another substantial limitation is the necessity for users to provide detailed text prompts - a task in graphic design generation that typically requires a high degree of professional expertise to develop effective prompts. In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating superior quality graphic design images while only necessitating simple user intention.
We maintain that these three limitations critically undermine the quality of the generated graphic design images. A scalable, high-quality graphic design generation system should ideally require minimal effort from users, produce accurate and high-quality typography information for a variety of purposes, and offer a flexible editing space. This would allow users to further refine the output, integrating human expertise when necessary. Our objective in this work is to establish an efficient and dependable autonomous text-to-design system capable of generating high-quality graphic design images from user intent prompts. We suggest breaking down the complex process of graphic design image generation into a hierarchical generation process. This process involves multiple specialized generation models, each designed to handle different sub-tasks:
First, the focus is on creative design and interpretation, essentially focusing on understanding intentions. This is accomplished by employing state-of-the-art LLMs, specifically the Llama-B [33] and fine-tuning it with a comprehensive dataset of nearly curated intention-JSON pairs. The JSON file contains critical design-related data, such as background captions, object captions, and textual details. Additionally, we provide optional specifications for object positioning and others.
Second, our focus shifts to visual composition and enhancement, encompassing two sub-tasks: the generation of visual elements and the creation of typography attributes. This involves fine-tuning specialized cascaded diffusion models like DeepFloyd/IF [19] to generate diverse visual elements. The construction of these models ensures the harmonious integration of elements like the embellished background and object image layers. We then use a typography Large Multimodal Model (LMM), built with LLaVA--B [24], to predict the typography JSON file. This leverages the Design LLM’s forecasted JSON file, a diffusion model’s projected background image, and a cascaded diffusion model’s anticipated object image. A graphic renderer then assembles these elements according to the layout in the predicted JSON file.
Third, the process concludes with feedback and quality assurance, aimed at enhancing the overall design quality. This involves meticulously fine-tuning a reflection LMM and employing GPT-4V(ision) for a thorough, multi-faceted quality analysis. This final step facilitates the necessary refinement of the JSON file, including adjustments to the text box position and sizes as deemed necessary.
Last, to access the capability of our system, we construct a DesignerIntention, which consists of professional graphic design intention prompts covering diverse categories and creative ones, to compare our method with existing state-of-the-art image generation system, conduct thorough ablation experiments for each generation model on different sub-tasks, offer a detailed analysis of the graphic designs created by our system, and engage in a discussion about both the limitations and prospective future pathways of graphic design image generation.
2 Related Work
Autonomous Agent based on LLMs Recent studies have proposed the development of advanced agent systems leveraging powerful Large Language Models (LLMs) to tackle a variety of challenging tasks. We can classify these into two primary categories: (i) General Multi-Agent Systems: These attempts aim to create an agent capable of addressing as many tasks as possible, ranging from coding, mathematics, science to societal related tasks. Among these attempts, several studies [11, 27, 40, 6, 40, 22] have explored the similar communicative (or cooperative) multi-agent framework to deconstruct a complex task into multiple sub-tasks, each handled by a specialized agent. For instance, ChatDev [27] introduces a chat-based software development framework to break down the development process into sequential atomic subtasks, where each subtask demands collaborative interaction and cross-examination between two roles. (ii) Image Generation Agents: These studies [45, 20, 39, 44, 29] concentrate more on image generation. For example, the concurrent work Idea2Img [45] unveils a framework for automated image generation, supporting complex prompts, and enhancing any text-to-image model for image design. Mini DALLE3 [20], a DALLE3 derivative, provides an interactive text-to-image generation system for a seamless user experience.
Our model primarily distinguishes itself from existing image generation agent studies [45, 20] and multi-agent efforts [27, 40, 22] in the mechanism of task decomposition and the usage of tools. While most existing studies incorporate tool usage knowledge by designing tailored prompt instructions or in-context examples for GPT-/, or GPT-4V(ision) to iteratively refine image generation prompts, we empirically demonstrate that these approaches fall short in excelling at professional graphic design image generation tasks due to the complexity of the required expertise. Unlike these efforts, our Cole decomposes the complicated graphic design image generation task into a hierarchical set of generation tasks, and performs dedicated fine-tuning of each generation model on our constructed professional structured graphic design dataset.
Design Image Generation Graphic design generation, given its extensive applications in advertising, branding, and marketing, is a vital research area with substantial industry impact. As discussed in [13, 9, 12, 21, 8, 23, 4, 46], a well-crafted graphic design can effectively convey messages through a skillful blend of shapes, colors, and fonts, all while respecting design principles and aesthetics. A significant amount of research has been invested in generating specific components of professional graphic design images, such as layouts or typography attributes. For instance, FlexDM [16, 43] leverages a unified Transformer architecture for element prediction. Additional research [30] focuses on typography generation, while UI-Diffuser [35] investigates mobile UI generation. On the other hand, the majority of studies [12, 23, 4, 46] primarily concentrate on layout generation, assuming that all other information is at hand.
Unlike the above-mentioned methods that focus on individual aspects within the graphic design generation task, this study aims to develop an autonomous, comprehensive system capable of executing all design generation tasks from clear user intent descriptions. This task formulation presents a significantly greater challenge compared to most existing approaches. To the best of our knowledge, our Cole is among the pioneering works that seek to tackle the complex task of autonomous graphic design generation based solely on a basic user intent prompt.
3 Our Approach
3.1 Cole Framework Overview
Motivated by the necessity to generate high-quality graphic designs that accurately reflect user intentions, we recognize the importance of a seamless integration of multiple roles. Each of these roles is designed to address specific aspects of the graphic design process, as indicated in the references [36, 34]. Consequently, we have broken down the text-to-design generation task into a hierarchical structure of tasks. These consist of the intention-to-JSON generation task, text-to-background image generation task, text-to-object image and alpha mask generation task, typography attribute generation task, and the graphic rendering task. The tasks of intention-to-JSON generation and typography attribute generation both follow a hierarchical structure. They necessitate the prediction of missing values given a predefined structured request.
We introduce the comprehensive hierarchical framework of our Cole system as depicted in Figure 3. To effectively train the specialized design-agent system, we initially amass approximately 100,000 high-quality raw graphic design images from the internet. Following this, we organize these graphic design data into a hierarchical structure which includes components such as (intention, JSON) pairs, (caption, background-image) pairs, (caption, object-image) pairs, and (text-image, typography JSON) pairs. We will further explore the process of constructing each component in the ensuing discussion.
3.2 Design LLM
We first formulate the intention-to-JSON prediction task as a masked field prediction problem following [15]. This approach learns to predict element types, positions, attributes, and text using a unified Transformer architecture. Furthermore, we adopt a strategy from [1] to construct causally masked multi-modal LLMs by leveraging the Llama-B [33].
We have fine-tuned the causally masked Llama-B model using our assembled dataset of nearly (intention, JSON) pairs, over an approximate span of epochs. Our empirical findings suggest that incorporating additional predictions of the complete caption, along with key words from the graphic design image, facilitates subsequent tasks of background and object image generation.
Given that both intentions and detailed captions are absent in the raw graphic design data, we first employ GPT- and GPT- models to generate the intentions. This is accomplished with the aid of the system prompt displayed in the “Intention Generation Prompt”. For a more comprehensive understanding, we have included detailed intention generation prompt with in-context examples in the supplementary materials. Subsequently, we utilize the recently developed, large multimodal model known as LLaVA--B model [24]. This model is effective in generating detailed captions for the entire image, background image, and object image captions.
Figure 4 presents a representative example, showcasing the ground-truth triplets of graphic design images, alongside the generated user intentions and corresponding JSON files. In addition, we have also investigated the potential of the highly proficient GPT- model for directly generating JSON from user intentions. This is achieved by formulating an instructive prompt. Our empirical analysis, detailed in the experiment section, demonstrates that our methodology consistently surpasses the performance of GPT-.
3.3 Text-to-Background Diffusion Model
Given the current unavailability of mature text-to-SVG technology and the scarcity of high-quality text-SVG pairs, which leads to the absence of strong text-to-SVG generation models [17, 47, 41, 42], our study suggests consolidating all SVG elements and additional embellishments into one unified image layer. Therefore, we first extract the generated detailed description of the background image layer from the JSON file. Then we train a text-to-background diffusion model to generate the coherent background images consisting of all the SVG files and decoration elements.
We opt to fine-tune our collected less than (caption, background image) pairs using the recent DeepFloyd IF [19] diffusion models for epochs. The output resolution is . Following this, we employ the generated background images as conditioning input for the text-to-object diffusion model, which predicts coherent high-resolution object image layers based on their corresponding object prompts.
3.4 Text-to-Object Diffusion Model
Subsequently, we propose training a text-to-object diffusion model capable of perceiving both the object captions present in the JSON file and the visual pixels predicted by the text-to-background diffusion models.
Following the instruct-P2P approach [3], we concatenate the noise latent vectors with the latent vectors of the conditioning background image with the alpha mask for the object layer, and we add additional input channels to the first convolutional layer. All available weights of the diffusion model are initialized from pre-trained checkpoints, while weights operating on the newly added input channels are initialized to zero. As a result, we can predict both the object image and its alpha mask layer, which enables us to control the transparency of the object pixels. Consequently, we can seamlessly blend the object image layer with the background image layer in accordance with the alpha mask. Empirically, we found that predicting the additional composed image is pivotal to ensuring a more reasonable placement of the object pixels. Therefore, our text-to-object diffusion model needs to predict channels. For more details, please refer to the experimental section.
Similar to the text-to-background generation model, we optimize our assembled dataset, comprising approximately caption and object image pairs. The relatively smaller number of pairs is due to the frequent absence of the object image layer in many graphic design images. We employ the instruct-P2P scheme for optimization, which is based on the DeepFloyd [19] IF diffusion models, and this process is carried out over an estimated span of epochs. Figure 5 provides a selection of representative examples that demonstrate our method of decomposing complex graphic design images into distinct layers of background and object images. It’s important to note that predicting the object alpha mask is essential for facilitating flexible editing operations on the object pixels. However, ensuring accurate alpha mask prediction presents a complex challenge. Please refer to the supplementary materials for more in-depth technical details.
3.5 Typography LMM
Typography [38] is the art and technique of arranging type to make written language legible, clear, and visually engaging for the reader. More succinctly, typography refers to the style and appearance of text. It encompasses the design of font style, appearance, and structure, with the goal of eliciting specific emotions and conveying certain messages. In essence, typography brings text to life.
Contrary to recent efforts, such as Tanveer et al. [32], Iluz et al. [14], and He et al. [10], which primarily employ diffusion models to generate novel letter-level or word-level semantic or artistic typographies, our focus is more directed towards the utilization of an LMM (Large Multimodal Model). This approach allows us to address the highly challenging task of image-level typography. This task involves the selection of font types, establishment of hierarchy arrangements, adjustment of leadings (line spacing), tracking (character spacing), and kernings (space between specific characters).
In our experiments, we elected to fine-tune the LLaVA--B model [24] using nearly triplets for epochs. These triplets consist of multiple texts, composed images of backgrounds and objects, and typography information. Figure 6 displays a representative example of a Typography JSON file. It is evident that predicting visually appealing typography information for each text box, based on the input image, is a complex task. This is due to the requirement of predicting approximately attributes for each box. To our knowledge, even the most recent effort [30] only manages to predict typography attributes, assuming the other attributes are fixed. To simplify the prediction process, we further categorize the continuous floating numbers into discrete bins as per the method outlined in [5] and we also utilize a unique token to represent the values that fall within the same bin. We explain more details of the bin partition scheme in the supplementary material.
3.6 Graphic Render
As demonstrated in the preceding sections, we have successfully predicted the JSON file, complete with detailed captions and text elements. We have also generated the background and object image layers (with object alpha mask), ensuring they adhere to the principles of design aesthetics and layout. In addition, we have produced the typography JSON file, which precisely outlines the style and appearance of the text elements.
We implement a graphic renderer founded on Skia [31], an open-source 2D graphics library. It is important to highlight that the rendering quality, as per the official implementation of CanvasVAE [43], presents significant potential for enhancement. Figure 7 showcases a qualitative comparison of the results obtained from our rendering implementation, contrasted with those generated by CanvasVAE. Upon evaluation, it is evident that our graphic renderer significantly outperforms the open-sourced CanvasVAE.
Original
Ours
CanvasVAE
Original
Ours
CanvasVAE
3.7 Reflect LMM & Quality LMM
We find that initial graphic design outputs may not always align with user expectations. Therefore, we propose the integration of an added Reflect-LMM. This component is designed to critically analyze the generated design images and predict necessary adjustments to the JSON file, thereby further improving the resulting graphic designs.
The entire self-reflection framework is illustrated on the left side of Figure 8. As an integral part of our implementation plan, we fine-tune a LLaVA--B model on a dataset that comprises more than artificially generated noisy JSON files. Each of these files is paired with a corresponding ground-truth JSON file. Our primary modification involves the addition of random shift noise to the position of each ground-truth typography block.
Besides, we utilize GPT-4V(ision) [25] to evaluate the quality of the generated graphic design images, as displayed in the “Quality Assurance Prompt for GPT-4V(ision)” in the appendix. Through empirical analysis, we have found that GPT-4V(ision) excels at evaluating the quality of main aspects such as design and layout, content relevance, graphics and images, and innovation. However, it falls short in assessing the quality of typography and color. In addition, we also incorporate the evaluation prompt design as introduced in a recent concurrent effort [18]. This evaluates the graphic design images from approximately aspects, including text rendering quality, composition and layout, color harmony, cinematography, style, image-text alignment, aesthetics, and overall design.
4 Experiment
(a) of text box, words, chars
(b) Wordcloud
| Category |
Intention Examples |
| advertising | Create a customer testimonial advertisement showcasing a pink clothing collection. The advertisement includes a positive review from Jenny Wilson. |
| events | Create a birthday card with a festive and colorful design featuring a cake with fruits. The card is an invitation to a birthday party happening on June 22nd from 8pm to 11pm at 123 st. Any city 4567. |
| marketing | Design a gift certificate for a travel insurance company that offers a 30% discount for all kinds of travel. |
| posts | Create an Instagram story to promote a special offer of 30% discount on the autumn collection. Encourage users to visit the website by including a link in bio. |
| covers& headers | Create a thumbnail for a vlogger episode that features a collection of rare antique telephones. The title is ’Top 5 Antique Phones’ |
| creative | Design an alluring poster for a blues concert in a clearing of the Amazon Rainforest. The artwork should evoke the soulful tunes of blues music echoing through the wilderness. Use a moody, midnight blue palette and incorporate silhouettes of blues instruments and nocturnal wildlife. |
(a) covers & headers
(b) marketing
(c) posts
(d) events
(e) advertising
background image
object image
object alpha mask
composed image
To assess our Cole’s proficiency in generating graphic design images, we first build a DesignerIntention benchmark comprising over approximately graphic design intention prompts. These prompts serve as a tool to evaluate the model’s performance from diverse perspectives. Second, as mentioned in the previous section, we propose utilizing GPT-4V(ision) to evaluate the quality of the generated graphic design images and compare our system with the latest systems including DeepFloyd/IF(+GPT-), SDXL(+GPT-), and DALLE3 (+GPT-) following [2, 18]. Last, we carry out ablation experiments to scrutinize the performance of each component and discuss the limitations of our system.
4.1 DesignerIntention Benchmark
We have constructed a benchmark for designer intentions, amassing approximately intention prompts for each of the most pertinent categories, namely marketing, covers and headers, posts, events, and advertising. In a complementary effort, we have compiled an extra set of around unique, creative user intentions. The graphic design images derived from these inventive prompts are displayed for your reference in last row of Figure 16. Refer to the supplementary for an all-inclusive review of these roughly intention prompts, along with the creative ones. Figure 9 furnishes comprehensive statistics of our benchmark from myriad perspectives. We conduct all the experiments on DesignerIntention by default.
4.2 Main Results
Comparison with State-of-the-art
Figure 15 presents a comparison with various state-of-the-art text-to-image generation models that are augmented with GPT-. The usage of GPT- stems from the fact that traditional text-to-image models generally necessitate detailed descriptions, which are often missing in the intention prompt. Consequently, we utilize GPT- across all models to translate the user’s simple intentions into intricate captions. The generation results are then evaluated accordingly across different systems. We provide the original intention prompts in Table 1. We provide a detailed enhanced prompts, augmented using GPT-, within the supplementary material.
In addition, we showcase their predicted numerical quality scores from GPT-4V(ision) based on quality assurance prompt in Figure 10. A careful evaluation reveals that our Cole achieves very competitive quality accessing results even compared to the latest DALLE3. Our Cole achieves the best quality when generating covers & headers or posters. An interesting observation is that the background image generated by DALLE3 is of exceptional quality but the text regions are non-editable. We also empirically show some impressive results of our Cole when replacing our text-to-background generation model with DALLE3 by leveraging the text segmentation models, inpainting models, and others.
Seed #1
Seed #2
Seed #3
Seed #4
Seed #5
initial image
+ object editing
+ text editing
+ combined editing
Ours
DeepFloyd/IF
SDXL
DALLE3
Qualitative results
Figure 16 showcases a variety of graphic design images entirely created by our system. It’s noteworthy that our system performs exceptionally well, considering that each model within it is trained independently in a hierarchically organized manner. However, we recognize several limitations within our Cole system, including: (i) the arrangement of the typography block, (ii) the limited number of editable visual elements in the image, and (iii) the restricted diversity in typography color selection. Addressing these issues is a direction we’d like to pursue in our future work.
4.3 Ablation Experiments
Design LLM
We assess the performance of the Design LLM by calculating the semantic textual similarity (STS) between the predicted text and the ground-truth ones using Sentence-BERT [28].
This evaluation is done across various keys, such as global or background or object caption, keywords, and heading&sub-heading&bodytexts, following the generated JSON file. Their average STS scores are , , , , and , respectively.
Text-to-Background & Text-to-Object
Figure 12 showcases the generated background image layer, object image layer, and the composed image layer. The displayed visualizations highlight the proficiency of our text-to-object generation model in crafting a coherent object image layer and its corresponding alpha mask. We have also carried out a detailed comparison of the background images generated by our model and others [26, 19, 2]. Refer to the supplementary material for more details.
Typography LMM & Reflect LMM
We visualize the predicted results with the ground-truth images as the condition in Figure 12. We assess these results from two perspectives: typography scoring with GPT-4V(ision) and the localization accuracy including , , and . For a more comprehensive comparison and the ablation of the reflect-LMM, refer to the supplementary material.
Design Diversity
Designers often face the critical challenge of delivering diverse designs while maintaining high quality. Our Cole system empirically excels in this regard, as evidenced by its operation with various random seeds shown in Figure 14.
Flexible Editing
Our Cole offers a versatile and interactive platform that allows users to edit both the object image layer and typography. This functionality augments the capability to manipulate generated graphic design images from various perspectives. Figure 14 showcases examples of different editing operations. For instance, users have the flexibility to resize, reposition, or replace the object image layer, as well as to freely modify the font type and size.
5 Conclusion
This paper introduces an exceptionally efficient hierarchical generation framework., Cole, that simplifies the intricate process of graphic design generation by breaking it down into manageable, coordinated sub-tasks, each yielding distinct outputs. Leveraging the benefits of this hierarchical task decomposition, our system can generate competitive graphic design images from simple user intent even when compared to DALLE3 on the challenging DesignerIntention benchmark. The core strengths of our Cole system lie in ensuring precise text rendering and providing flexible editing options. We have also pinpointed several key limitations of our Cole system in the context of generating graphic design images, and we plan to tackle these in future research.
Acknowledgement We extend our gratitude to Xinyang Li for his invaluable assistance with data preparation. Our thanks also go to Yang Ou for providing insightful comments from a professional designer’s perspective. We appreciate the enlightening discussions offered by Chunyu Wang, Jianmin Bao, Mingxi Cheng, and Chong Luo. Additionally, we want to give a special mention to Weicong Liang for his indispensable help in addressing the LaTeX compilation errors.
Supplementary Material
A Graphic Design Pipeline
The generation of high-quality graphic designs from user intentions requires the harmonious integration of several roles or components, each tailored to address specific aspects of the graphic design process [36, 34]. We initially offer a detailed plan that outlines these roles, clarifies their purposes, and explains their interactions. Figure 17 illustrates the complete pipeline.
Creative Design and Interpretation
The primary function of this role involves understanding user intentions and translating them into tangible design elements like titles, subtitles, headings, subheadings, and body text. It also sets the tone and style of the graphic design, which incorporates elements such as primary color, theme, style, and captions.
Visual Composition and Enhancement
The key responsibility of this role is to uphold the principles of design aesthetics and layout, which involves the generation, assessment, and adjustment of visual appeal and spatial organization. The generation process mainly utilizes a comprehensive library of images, icons, and textures to enhance the design. This module further refines the initial concepts originating from the Creative Design and Interpretation engine, optimizing the layout and enriching the design with suitable assets. For instance, it is imperative to not only generate images of objects, decorations, and backgrounds, but also to predict various typographic attributes for elements such as the heading, subheading, and body text.
Feedback and Quality Assurance System
The principal function of this role is to integrate user feedback into the design in an iterative manner, with the objective of ensuring that generated designs comply with quality benchmarks and regulatory standards. This module guarantees that the designs not only align with user expectations, but also fulfill professional and legal criteria. It maintains a continuous interaction with the first two modules to guide the design refinement process.
Collaboration Loop
This system ensures that the designs not only meet user expectations, but also adhere to professional and legal standards. It maintains a constant dialogue with the first two modules to navigate the design refinement process effectively.
Discussion
Unlike traditional graphic design systems that rely significantly on the expertise of professional designers, we aspire to create a DesignerAgent. This innovative system is designed to generate superior graphic designs, leveraging user intention descriptions alone. By fine-tuning state-of-the-art Large Multimodal Models (LMMs), Large Language Models (LLMs), and diffusion models, we aim to fully exploit our carefully curated, high-quality graphic design datasets to achieve this objective.
B. Detailed System Prompts for GPT-3.5 and GPT-4V(ision)
We introduce the Intention Generation Prompt for GPT3.5, Quality Assurance Prompt for GPT-4V(ision), and Pair-wise Comparison Prompt for GPT-4V(ision) as follows: The Intention Generation Prompt is crafted to deduce the user’s purpose utilizing diverse image data, such as keywords, all text visible in the image, the image’s title, among others. The Quality Assurance Prompt for GPT-4V(ision) is established to evaluate the quality of the created graphic design images across roughly five dimensions. These dimensions include design and layout, content pertinence, graphics and imagery, creativity, and the use of typography and color. The Pair-wise Comparison Prompt for GPT-4V(ision) is proposed to objectively compare our generated images with results from other methods from six perspectives of graphic design.
C. The Detailed List of DesignerIntention
A comprehensive list of designer intention prompts is summarized across Table 4, Table 4, Table 4, and Table 6.
Table 4 details the intention prompts corresponding to the images exhibited in Figure 1 of the main paper. In Figure 20, we show a few example pairs that illustrate the original image information alongside the corresponding generated intention prompt.
Table 4 presents the intention prompts augmented with GPT-, which were utilized for DALLE3 in Figure 2 of the main paper.
In Table 4, we display the augmented intention prompts with GPT- that were employed for DeepFloyd/IF, SDXL, and DALLE3 in Figure 10 of the main paper.
Last, Table 6 provides the detailed intention prompts for the design images generated, as depicted in Figure 11 of the main paper.
D. More Ablation Experiments
Text-to-Background & Text-to-Object To assess the efficacy of our text-to-background diffusion model, we conducted a comparative analysis involving DeepFolyd/IF, SDXL, and DALLE3, as depicted in Figure 21. The complete list of prompts for generating corresponding background images is listed in Table 6. Our findings reveal a notable trend among current state-of-the-art diffusion models: they often generate objects that occupy the entirety of the image, neglecting adequate space for accommodating visual text.
| Category |
Intention Prompt |
| creative | Design an animated digital poster for a fantasy concert where Joe Hisaishi performs on a grand piano on a moonlit beach. The scene should be magical, with bioluminescent waves and a starry sky. The poster should invite viewers to an exclusive virtual event. |
| advertising | Create a customer testimonial advertisement showcasing a pink clothing collection. The advertisement includes a positive review from Jenny Wilson. |
| marketing | Create a poster advertising a corner spice shop with a wide range of spices and ingredients. |
| posts | Promote a stand-up comedy show happening on November 12th, 20xx. |
| marketing | Design a gift certificate for a travel insurance company promoting their services with a 30% discount offer. The certificate should convey a sense of safety and protection while travelling. |
| posts | Design a recruiting post for a graphic designer job opening. Encourage interested candidates to send their CV to yourname@email.com |
| marketing | Design a Christmas sale poster with a fun and festive look. The poster should advertise up to 50% off on Christmas accessories, with a limited time offer. |
| marketing | Design a gift certificate for a travel insurance company that offers a 30% discount for all kinds of travel. |
| creative | Design an alluring poster for a blues concert in a clearing of the Amazon Rainforest. The artwork should evoke the soulful tunes of blues music echoing through the wilderness. Use a moody, midnight blue palette and incorporate silhouettes of blues instruments and nocturnal wildlife. |
| posts | Create a LinkedIn post with a cartoon illustration to promote an article about the benefits and growing popularity of remote work from home. The post links to a website for more information about the topic and is associated with a specific company and CEO John Doe. |
| marketing | Create a poster advertising babysitting services. The design should be warm and inviting, featuring a little girl and emphasizing professional and reliable childcare. The business name and phone number should be prominently displayed (+123 456 7890). |
| marketing | Create a poster advertisement for a restaurant’s happy hour special, showcasing the luxury and elegance of the establishment. The poster should include the restaurant’s name, location, website, and phone number for reservations and delivery. Happy Hour runs from 7 PM to 11 PM. |
| marketing | Create a real estate rental property flyer with a cozy and inviting interior design. Showcase the apartment’s features, amenities, and price. Include contact information for the real estate agency and their website address. The monthly rental price is $13,950. Bedrooms: 2, Bathrooms: 3, Square Footage: 1500. |
| creative | Design an interactive digital invitation for a space-themed gala. The invitation should feature an animated rocket launch sequence and include a scannable QR code that plays a personalized greeting from a holographic astronaut. |
| creative | Create a series of posters for a fictional Ancient Greek Olympics, with each poster highlighting a different sport like chariot racing or discus throw, styled in traditional Greek pottery art. |
| covers& headers | Create an Instagram story promoting a special Christmas offer for a chocolate drink. The offer includes a 20% discount and encourages customers to order now. |
| Category |
Intention Prompt |
| posts | An eye-catching car sale ad featuring a glossy white hatchback automobile, gleaming under the glow of a soft spotlight. The car’s sleek contours and sparkling rims are accentuated. Bold, crisp letters in the corner proclaim the enticing price of $13,900, with a monthly payment estimate of $179 highlighted beneath. A smart, clickable URL, CALIPSUN.CO.UK, is displayed prominently for further details, serving as a beacon for potential buyers. |
| posts | A grand display of our exclusive timepiece collection, each watch exuding elegance and style, set against a sophisticated black background. The watches are embedded with intricate details and shimmering dials. A bold, eye-catching ’40% off’ sign is placed prominently, urging viewers to upgrade their style. On the top corner, the Instagram Story’s ’Swipe Up’ prompt is nestled, symbolizing the immediacy of this limited-time offer. |
| advertising | A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged. In the corner, a gleaming starburst accentuates a glowing testimonial from Jenny Wilson, her words of praise shimmering. Beneath the testimonial, elegantly scripted contact information is highlighted: phone number 123-456-789, website WWW.FASHIONSTYLE.COM, and an address at 123 Anywhere ST., Any City. |
| posts | An energetic poster with a vibrant color palette, showing joyful children safely riding top-quality scooters. Bold text announces, ”Back to School Scooter Sale!” A prominent ”40% off” starburst highlights the best scooters for kids. Parents are seen looking on contentedly, knowing they’ve chosen a fun, safe ride for their children to school. |
| Category |
GPT4 Refined Intention Prompt |
| advertising | A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged. In the corner, a gleaming starburst accentuates a glowing testimonial from Jenny Wilson, her words of praise shimmering. Beneath the testimonial, elegantly scripted contact information is highlighted: phone number 123-456-789, website WWW.FASHIONSTYLE.COM, and an address at 123 Anywhere ST., Any City. |
| events | A vibrant birthday card bathed in festive hues of red, orange, and yellow. A delectable cake, crowned with an assortment of juicy fruits, serves as the focal point. Around it, confetti and streamers tumble playfully. The card invites you to a birthday celebration on June 22nd, from 8pm to 11pm, in a quaint house located at 123 Street, Any City 4567. |
| marketing | Picture a vibrant gift certificate for a leading travel insurance company. A globe subtly watermarked in the background, planes, and ships creatively darting across. Emblazoned on the center in bold, gold lettering is the irresistible offer: ”30% Discount on All Travel Insurance Policies”. Borders of the certificate are adorned with famous landmarks, adding a touch of wanderlust. The company’s logo sits proudly at the top, promising security and peace of mind for all journeys. |
| posts | A vibrant autumnal landscape with leaves in shades of gold, amber, and crimson cascading from the trees. A sleek banner sweeps across the scene, displaying the text ”30% off Autumn Collection”. A stylish arrow gently points downwards towards a glowing ”Link in Bio” button. The whole image exudes an irresistible allure, beckoning Instagram viewers to partake in this exclusive offer. |
| covers& headers | A vintage-inspired thumbnail for a vlogger episode, prominently featuring a curated assortment of five rare antique telephones, each unique in design and era. The collection is artfully arranged on a mahogany table against a backdrop of antiquated maps and time-worn books. The eye-catching title, ’Top 5 Antique Phones’, is elegantly superimposed on the scene in a classic, bold typeface. |
| creative | Imagine a captivating poster for a blues concert nestled in a moonlit clearing of the Amazon Rainforest. The midnight blue canvas paints an enchanting scene of silhouetted blues instruments - a saxophone, a guitar, and a harmonica - intertwined with the silhouettes of nocturnal wildlife like sleek panthers and hooting owls. The artwork pulsates with the soulful rhythm of blues music, echoing through the wilderness, stirring the soul. |
| Category |
Intention Prompt |
| covers& headers | Create a Youtube thumbnail for a video showcasing how to make Italian dishes at home, specifically focusing on rolling out dough. |
| covers& headers | Create a Youtube thumbnail to promote a pastry making tutorial series. The series is called ’The Best Pastry in the World’ and is hosted on the channel ’Your Channel’. The thumbnail should showcase a collage of homemade pastries, specifically croissants. Part 01 should be emphasized. |
| covers& headers | Create a Youtube thumbnail for promoting an antique store subscription service. Showcase the variety of unique vintage goods available in the store. |
| covers& headers | Design a Facebook event cover for a winter decoration workshop featuring lectures and demonstrations by Miranda Yates, Sam Knightly, Zachary Evans, and Veronica Floyd on December 24 at Lafayette Studio. |
| covers& headers | Create a Facebook cover image to announce a flower exhibition event. |
| marketing | Create a Thanksgiving gift certificate with an elegant and sophisticated design that offers a 50% discount on decorations. The message wishes a happy Thanksgiving to the recipient. |
| marketing | Design a colorful store logo with an illustration of leaves to represent a fresh and vibrant shopping experience. The store may be named TB. |
| marketing | Create a coupon that promotes a garden design service offered by Yard Culture. The coupon is a gift voucher valued at $49 and can only be used once. It is valid until November 15th. The coupon number is No:7496-4587. |
| marketing | Design a poster to promote a charity bike ride event to raise funds and support for a humanitarian cause. The event will be held on June 26th and the website for registration and donation is www.funbike.com. |
| marketing | Create an advertisement poster for a college merchandise store that has been operating since 1991. The design should focus on college and university branded gear, including clothes, backpacks, and bags. |
| posts | Create an Instagram story frame with a futuristic and fantasy-like design that features a man in an astronaut helmet and a space-themed background. The text #ThisIsMetaLife and SPACEDOUT are included. |
| posts | Create a recruiting Instagram post to advertise a job vacancy for a UI/UX designer. Encourage passionate individuals to apply and provide a link to apply in the bio. |
| posts | Promote a personal styling service for summer vacation outfits on Instagram. The contact number is +00 1357 2468 and website is www.website.com. |
| posts | Create an Instagram story promoting a special Christmas offer for a chocolate drink. The offer includes a 20% discount and encourages customers to order now. |
| posts | Create an Instagram story ad for an antique shop, promoting the sale of a timeless typewriter. Encourage customers to visit the store by swiping up on the story. |
| events | Design a holiday-themed gift card to promote the sale of gift boxes and packaging for the Christmas in July event. Include the website address www.yourbrand.com and store name. |
| events | Create a Valentine’s Day greeting card with a cute and romantic design. The card includes the phrase ’Happy Valentine’s Day’ and ’Much love to you!’ |
| events | Create a business card for a car service company that highlights their commitment to providing reliable and efficient car maintenance and repair services. The slogan ’Give your car wings’ emphasizes the quality of service provided. |
| events | Design an advertisement card to promote the sale of premium quality school equipment, with a focus on enhancing the learning experience for students. |
| events | Create a Women’s Day greeting card with signs of female gender, celebrating women and their achievements. The design should be beautiful and include floral elements. The card should be suitable for store use and have a message of ’Happy Women’s Day’. |
| advertising | Create an Instagram ad for a real estate agency highlighting the modern and prestigious skyscraper building they have on offer. Encourage viewers to ’reach new highs’ by visiting the website at www.togalim.eu. Include contact phone numbers 126-3764-1376 and 301-17361-5085 for interested clients. |
| advertising | Create an Instagram ad for a real estate agency highlighting the modern and prestigious skyscraper building they have on offer. Encourage viewers to ’reach new highs’ by visiting the website at www.togalim.eu. Include contact phone numbers 126-3764-1376 and 301-17361-5085 for interested clients. |
| advertising | Create a Facebook advertisement for a store selling natural essential oils that promote healing for the body and soul. Highlight the use of floral and plant ingredients in the product. Encourage viewers to visit the store. |
| advertising | Design a weekend sale advertisement for skincare products with a focus on promoting the cosmetic serum. All items are discounted up to 30% for this season. |
| advertising | Create an Instagram ad for a camping festival and hiking event. The ad should promote the event and provide information about registration, including a website. Additionally, offer tips for surviving while hiking. |
| advertising | Create an Instagram ad for a toy store’s Black Friday and Cyber Monday sale with a focus on the discount of 45% off on selected toys. |
| creative | Design an attractive poster for a Mars botanical garden tour. Illustrate a greenhouse with a variety of extraterrestrial plants. Use vibrant colors to depict the potential biodiversity on Mars. |
| creative | Create an enticing poster for a classical orchestra concert in the heart of the Amazon Rainforest. Depict the harmony between the music and the wilderness with a blend of vibrant greens and subtle sepia tones. Incorporate subtle hints of wildlife and flora within the musical instruments. |
| creative | Design a dramatic poster for an opera performance in a natural Amazon Rainforest amphitheater. Use a bold, Baroque-inspired style to reflect the grandeur of opera, juxtaposed with the raw beauty of the forest. Show a dramatic scene unfolding on the forest stage. |
| creative | Create a poster for an imaginary music festival where classical composers like Mozart and Beethoven are headliners. The design should merge classical elements with a modern music festival vibe, including stage setups and audience. |
| creative | Construct an eye-catching poster for an upcoming academic conference on Quantum Physics. The design should encapsulate the mysterious and profound nature of the subject. Use a black hole or quantum particles as the main design element, and apply a deep space color scheme to create an intriguing atmosphere. Make sure the conference details and keynote speakers are clearly visible. |
| Category |
Image Generate Prompt |
| marketing | Gift Certificate. blank, empty, layout, tb, certificate, offer, illustration, pattern, floral, flowers, pink, frame, sale. The image features a white background with a pink border. The border is adorned with a decorative floral pattern, giving the image a visually appealing and elegant appearance. The combination of the white background and the pink border with the floral pattern creates a harmonious and attractive design. |
| posts | Instagram Story. astronaut, cosmic, cosmonaut, helmet, orange, science, space, spaceman, universe, story, instagram, ig, insta, frame. The image features a man wearing a spacesuit, sitting on a planet with his legs crossed. He is positioned in the center of the scene, and the planet is displayed in the background. The man appears to be a spaceman, possibly an astronaut, as he is dressed in a suit designed for space travel. The scene is set against a dark blue sky, which adds to the overall atmosphere of the image. |
| marketing | Gift Certificate. agency, holiday, journey, service, tourism, travel, traveling, trip, vacation, tb, certificate, sale, discount, insurance, illustration, cartoon, blue. The image features a map with a variety of travel-related items scattered around it. There are several pairs of shoes, including a couple of boots, placed in different positions on the map. A hat is also present, positioned near the top left corner of the map. In addition to the shoes and hat, there are multiple books scattered across the map, with some located near the bottom left corner and others in the middle and top right areas. A pair of scissors can be seen in the top right corner of the map, and a clock is placed near the center of the map. |
| events | Business Card European. data, card, health, home, professional, service, technology, icon, house, nursing, care, modern, medical, health care, contact, information. The image features a house with a pink roof, which is situated next to a street. The house is surrounded by a white fence, and there is a person standing on the sidewalk in front of the house. The person appears to be looking at the house, possibly admiring its architecture or considering it for purchase. The scene also includes a car parked on the street, adding to the overall atmosphere of the neighborhood. |
| events | Instagram Post. wedding, dinner, rehearsal, banquet, event, holiday, celebration, marriage, food, party, restaurant, ceremony, leaf, jungle, tropical, exotic, brown, instagram, post, insta, ig, tb. The image features a brown background with a few green leaves scattered throughout the scene. The leaves are of various sizes and are positioned at different angles, creating a sense of depth and texture. The overall composition of the image is simple, with the brown background and green leaves being the main focus. |
| events | Instagram Post. empowerment, equality, feminism, holiday, march, march 8, rights, women, women’s day, womensday, tb, illustration, female, girls, celebration, beautiful, greeting, sisterhood, instagram, insta, ig, post, flowers, floral. The image features two people standing next to each other, each holding a heart. The hearts are positioned above their heads, creating a visually striking scene. The people are standing in front of a backdrop of flowers, adding a touch of color and beauty to the image. |
| covers& headers | Ticket DIN large. festival, art, abstract, geometric, graphic, illustration, pattern, shape, simple, visual, colorful, blue, ticket. The image features a colorful and abstract design, with a predominantly blue background. The design consists of various geometric shapes, including triangles, squares, and circles, arranged in a visually appealing manner. The shapes are positioned throughout the image, with some closer to the foreground and others further in the background. The overall composition creates a sense of depth and complexity, making the image visually engaging and interesting. |
| advertising | Square Graphic Post. tb, instagram, insta, ig, post, style, fashion, trendy, fashionable, stylish, beautiful, clothes, model, woman, posing, modern, female, young, casual, beauty, attractive, look, color, modeling, lady, people, outfit, order, new arrival, pink, customer, testimonial, review. The image features a pink background with a white square placed in the center. The square is surrounded by a pink hue, giving it a slightly pinkish tint. The background is predominantly pink, with no other objects or distractions visible in the scene. |
| Setting | Design LLM | Text-to-Background DM | Text-to-Object DM | Typography LMM | Reflect LMM | ||
| Stage 1 | Stage 2 | Stage 1 | Stage 2 | ||||
| Training epochs | |||||||
| Batch size | |||||||
| Learning rate | e- | e- | e- | e- | e- | e- | e- |
| AdamW | e- | e- | e- | e- | e- | e- | e- |
| AdamW | (, ) | (, ) | (, ) | (, ) | (, ) | (, ) | (, ) |
| Learning rate schedule | Constant | Cosine | Cosine | Cosine | Cosine | Cosine | Cosine |
| Warmup steps | |||||||
| Max token length | - | - | - | - | |||
| Training strategy | % Causal, % CM3 [1] | - | - | - | - | Causal | Causal |
| Tuned parameters | B | M | M | B | B | B | B |
| LoRA rank | - | - | - | ||||
| LoRA alpha | - | - | - | - | - | ||
| Input resolution | - | ||||||
As a result, the background images produced by our specialized diffusion models prove more conducive for integration with diverse elements, such as the object image layer and the typography image layer. This suitability arises from their ability to create backgrounds that allow for seamless incorporation of these various elements, fostering enhanced compositional flexibility.
Typography LMM & Reflect LMM In Figure 20, we report the localization accuracy comparison results, such as mIoU, AP, AP, and AP, for the predicted text boxes based on the Typography LMM and the refined text boxes derived from the Reflect LMM. To facilitate fair comparisons and to utilize the ground-truth annotations of the text boxes, we employ the ground-truth background image and the ground-truth text value as the visual and textual conditions, respectively. Following this, the combination of these ground-truth conditions is inputted into the Typography LMM to predict the various typography attributes, including the text box positions denoted by the values of left, top, width, and height. For the Reflect LMM, we input the rendered graphic design images (based on the Typography LMM) as the visual condition and the ground-truth text value as the textual condition, subsequently predicting the new values for left, top, width, and height.
To compute the mIoU, AP, AP, and AP scores, we initially compare each predicted text box with its corresponding ground-truth counterpart. Subsequently, these scores are averaged across all text boxes. As demonstrated in the comparison results in Figure 20, we observe a significant improvement in performance, particularly for the AP and AP metrics. For reference, we have illustrated the qualitative predictions derived from both the Typography LMM and Reflect LMM in Figure 20.
Comparison by Professional Designer and GPT-4V In order to conduct a thorough and precise evaluation of the image quality produced by our Cole and DALLE3 systems, we implement a strategy that incorporates both subjective and objective methodologies.
On the subjective front, we leverage the expertise of a professional graphic designer to appraise the images created by Cole and DALLE3. This analysis takes into account various factors such as composition, subjective impressions, and color scheme. An elaborate presentation of this assessment can be found in Figure 22. From an objective perspective, we create prompts akin to the Pair-wise Comparison Prompt for GPT-4V(ision), detailed in Section B, and input these prompts alongside two images into the GPT-4V(ision) model. We then conduct a comparative analysis of the images based on six distinct dimensions: overall composition layout, image composition, medium, overall coherence, overall aesthetics, and text rendering.
Specific outcomes of this comparison, along with the accompanying rationale, are illustrated in Figure 23. As evidenced by the two provided examples, our model consistently generates images that surpass those produced by DALLE3 in four key areas. This highlights our model’s superior performance, especially in terms of adherence to user intent and in the precision and readability of text rendering, as compared to DALLE3.
E. Hyperparameters
Table 7 outlines all the comprehensive training hyperparameters for our Cole system, which is composed of a hierarchical series of five generative models. The table is segmented into three parts: training settings, model settings, and data settings, listed in that order from top to bottom.
Typography LMM
+ Reflect LMM
Typography LMM
+ Reflect LMM
Ours
DeepFloyd/IF
SDXL
DALLE3

### 18 Mar 2024 v2

@@
-Cole: A Hierarchical Generation Framework for Graphic Design
+Cole: A Hierarchical Generation Framework for Multi-Layered and Editable Graphic Design
 Abstract
 Graphic design, which has been evolving since the 15th century, plays a crucial role in advertising.
-The creation of high-quality designs demands creativity, innovation, and lateral thinking.
-This intricate task involves understanding the objective, crafting visual elements such as the background, decoration, font, color, and shape, formulating diverse professional layouts, and adhering to fundamental visual design principles.
-In this paper, we introduce Cole 111In 1843, Henry Cole introduced the world’s first commercial Christmas card [37].
-We commemorate his contributions with the Cole system., a hierarchical generation framework purposed to comprehensively address these challenges.
-This Cole system can transform a straightforward intention prompt into a high-quality graphic design, while also supporting flexible editing based on user input.
+The creation of high-quality designs demands design-oriented planning, reasoning, and layer-wise generation.
+This intricate task involves understanding vague intentions and faithfully generating multi-layered visual elements, including the background, decorations, objects, and fonts, among others.
+It also requires layout planning for all elements and visual reasoning to generate all visual elements that satisfy visual design principles.
+Unlike the recent CanvaGPT, which integrates GPT-4 with existing design templates to build a custom GPT, this paper introduces the Cole 111In 1843, Henry Cole introduced the world’s first commercial Christmas card [36].
+We commemorate his contributions with the Cole system.
+system—a hierarchical generation framework designed to comprehensively address these challenges.
+This Cole system can transform a vague intention prompt into a high-quality multi-layered graphic design, while also supporting flexible editing based on user input.
 Examples of such input might include directives like “design a poster for Hisaishi’s concert.” The key insight is to dissect the complex task of text-to-design generation into a hierarchy of simpler sub-tasks, each addressed by specialized models working collaboratively.
@@
 Our hierarchical task decomposition can streamline the complex process and significantly enhance generation reliability.
-Our Cole system consists of multiple fine-tuned Large Language Models (LLMs), Large Multimodal Models (LMMs), and Diffusion Models (DMs), each specifically tailored for a design-aware text or image generation task.
-Furthermore, we construct the DesignerIntention benchmark to highlight the superiority of our Cole over existing methods in generating high-quality graphic designs from user intent.
-We perceive our Cole as an important step towards addressing more complex visual design generation tasks in the future.
+Our Cole system comprises multiple fine-tuned Large Language Models (LLMs), Large Multimodal Models (LMMs), and Diffusion Models (DMs), each specifically tailored for design-aware layer-wise captioning, layout planning, reasoning, and the task of generating images and text.
+Furthermore, we construct the DesignerIntention benchmark to demonstrate the superiority of our Cole system over existing methods in generating high-quality graphic designs from user intent.
+Last, we present a Canva-like multi-layered image editing tool to support flexible editing of the generated multi-layered graphic design images.
+We perceive our Cole system as an important step towards addressing more complex and multi-layered graphic design generation tasks in the future.
 1 Introduction
@@
 The pivotal factors contributing to these advancements include the use of powerful Large Language Model (LLM) as a text encoder, the scaling up of training datasets, increased model complexity, the design of advanced sampling strategies, and the enhancement of data quality, among others.
-We believe that now is the opportune moment to redirect our efforts towards more professional image generation, particularly in the area of graphic design generation considering its pivotal roles in advertising, branding, and marketing.
+We believe that now is the opportune moment to redirect our efforts towards more professional image generation, particularly the graphic design generation considering its pivotal roles in advertising, branding, and marketing.
 Graphic design [35], a professional discipline, leverages the power of visual communication to convey targeted messages to specific social groups, with clear objectives.
@@
 Another substantial limitation is the necessity for users to provide detailed text prompts - a task in graphic design generation that typically requires a high degree of professional expertise to develop effective prompts.
-In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating superior quality graphic design images while only necessitating simple user intention.
+In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating multi-layered and editable graphic design images from simple user intentions.
+The most recent study [17] has empirically shown that the groundbreaking DALLE3 exhibits impressive capabilities in generating high-quality design images, marked by visually striking graphics and layouts, as illustrated in Figure 2.
+However, these images are not without their flaws.
+They continue to confront several critical challenges, such as mis-rendered visual text, often resulting in missing or extra characters, a phenomenon also highlighted in [2].
+Furthermore, editing these generated images demands complex operations such as segmentation, erasing, inpainting, and others, owing to their inherently uneditable nature.
+Another substantial limitation is the necessity for users to provide detailed text prompts - a task in graphic design generation that typically requires a high degree of professional expertise to develop effective prompts.
+In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating multi-layered and editable graphic design images from simple user intentions.
 We maintain that these three limitations critically undermine the quality of the generated graphic design images.
@@
 This would allow users to further refine the output, integrating human expertise when necessary.
+| Method | Accuracy visual text | Multi-layered output | Flexible editing | Generated design | GPT4-V Score | ||||
+| Layout | Image | Typography | Content | Innovation | |||||
+| DeepFloyd/IF [18] | ✗ | ✗ | ✗ | ✓ | 65.31 | 56.59 | 55.32 | 83.69 | 78.23 |
+| SDXL [26] | ✗ | ✗ | ✗ | ✓ | 60.89 | 60.12 | 61.21 | 67.33 | 69.92 |
+| DALLE3 [2] | ✗ | ✗ | ✗ | ✓ | 80.29 | 69.77 | 74.73 | 89.16 | 90.61 |
+| CanvaGPT [4] | ✓ | ✓ | ✓ | ✗ | 75.66 | 57.01 | 75.35 | 70.10 | 63.27 |
+| Ours Cole | ✓ | ✓ | ✓ | ✓ | 80.82 | 72.53 | 75.46 | 87.44 | 88.72 |
 Our objective in this work is to establish an efficient and dependable autonomous text-to-design system capable of generating high-quality graphic design images from user intent prompts.
 We suggest breaking down the complex process of graphic design image generation into a hierarchical generation process.
-This process involves multiple specialized generation models, each designed to handle different sub-tasks:
-First, the focus is on creative design and interpretation, essentially focusing on understanding intentions.
-This is accomplished by employing state-of-the-art LLMs, specifically the Llama-B [33] and fine-tuning it with a comprehensive dataset of nearly curated intention-JSON pairs.
-The JSON file contains critical design-related data, such as background captions, object captions, and textual details.
-Additionally, we provide optional specifications for object positioning and others.
-Second, our focus shifts to visual composition and enhancement, encompassing two sub-tasks: the generation of visual elements and the creation of typography attributes.
-This involves fine-tuning specialized cascaded diffusion models like DeepFloyd/IF [19] to generate diverse visual elements.
-The construction of these models ensures the harmonious integration of elements like the embellished background and object image layers.
-We then use a typography Large Multimodal Model (LMM), built with LLaVA--B [24], to predict the typography JSON file.
-This leverages the Design LLM’s forecasted JSON file, a diffusion model’s projected background image, and a cascaded diffusion model’s anticipated object image.
-A graphic renderer then assembles these elements according to the layout in the predicted JSON file.
-Third, the process concludes with feedback and quality assurance, aimed at enhancing the overall design quality.
-This involves meticulously fine-tuning a reflection LMM and employing GPT-4V(ision) for a thorough, multi-faceted quality analysis.
-This final step facilitates the necessary refinement of the JSON file, including adjustments to the text box position and sizes as deemed necessary.
+This process involves multiple specialized generation models, each designed to handle different sub-tasks: First, we train a Design-LLM to comprehend vague intentions and undertake design-oriented planning tasks.
+We formulate this as a text-to-JSON prediction task, which requires transforming naive and sparse user intentions into structured, layer-wise captions or attributes essential for the following task decomposition.
+Second, to perform design-oriented visual reasoning and generation, we train additional three models including (i) a text-to-background diffusion model for background layer generation: it is crucial to guarantee that the generated background images reserve adequate space for the incorporation of various elements, which is different from conventional methods such as DALLE3, which often produce objects filling the entire image space.
+(ii) a text-to-object diffusion model for object layer generation, it’s essential to ensure the model to understand both the content and style of the background image.
+Subsequently, the model should generate an object layer that not only aligns stylistically but is also placed in logical positions.
+(iii) a Typography-LMM for typography prediction, it needs to predict distinct typography attributes for each text box.
+This component requires a sophisticated level of reasoning about the visual contents of the merged background and object image layers, then predicting the typography attributes accordingly.
+We also demonstrate the generalization capability of our Typography-LLM by applying it to processed design images originally generated using DALLE3.
+Refer to Figure 1 for the visual results.
+Besides, we also introduce a feedback and quality assurance process to enhance the quality.
 Last, to access the capability of our system, we construct a DesignerIntention, which consists of professional graphic design intention prompts covering diverse categories and creative ones, to compare our method with existing state-of-the-art image generation system, conduct thorough ablation experiments for each generation model on different sub-tasks, offer a detailed analysis of the graphic designs created by our system, and engage in a discussion about both the limitations and prospective future pathways of graphic design image generation.
+Figure 3 shows that, among both non-designers and designers, our Cole system outperforms DALLE3 in text fidelity and message conveyance, achieving win rates of 69.6% and 70.5%, respectively.
+Regarding visual appeal, non-designers show a slight preference for Cole, while designers favor DALLE3.
+Our results are also preferred by designers over those from the latest commercial product, CanvaGPT.
+Table 1 provides a detailed comparison highlighting the key differences between our approach, earlier text-to-image generation models, and the latest CanvaGPT.
+Additionally, we present GPT4-V scores from various perspectives, with our Cole outperforming others in three critical areas: layout, image, and typography quality.
 2 Related Work
-Autonomous Agent based on LLMs Recent studies have proposed the development of advanced agent systems leveraging powerful Large Language Models (LLMs) to tackle a variety of challenging tasks.
-We can classify these into two primary categories: (i) General Multi-Agent Systems: These attempts aim to create an agent capable of addressing as many tasks as possible, ranging from coding, mathematics, science to societal related tasks.
-Among these attempts, several studies [11, 27, 40, 6, 40, 22] have explored the similar communicative (or cooperative) multi-agent framework to deconstruct a complex task into multiple sub-tasks, each handled by a specialized agent.
-For instance, ChatDev [27] introduces a chat-based software development framework to break down the development process into sequential atomic subtasks, where each subtask demands collaborative interaction and cross-examination between two roles.
-(ii) Image Generation Agents: These studies [45, 20, 39, 44, 29] concentrate more on image generation.
-For example, the concurrent work Idea2Img [45] unveils a framework for automated image generation, supporting complex prompts, and enhancing any text-to-image model for image design.
-Mini DALLE3 [20], a DALLE3 derivative, provides an interactive text-to-image generation system for a seamless user experience.
-Our model primarily distinguishes itself from existing image generation agent studies [45, 20] and multi-agent efforts [27, 40, 22] in the mechanism of task decomposition and the usage of tools.
-While most existing studies incorporate tool usage knowledge by designing tailored prompt instructions or in-context examples for GPT-/, or GPT-4V(ision) to iteratively refine image generation prompts, we empirically demonstrate that these approaches fall short in excelling at professional graphic design image generation tasks due to the complexity of the required expertise.
-Unlike these efforts, our Cole decomposes the complicated graphic design image generation task into a hierarchical set of generation tasks, and performs dedicated fine-tuning of each generation model on our constructed professional structured graphic design dataset.
-Design Image Generation Graphic design generation, given its extensive applications in advertising, branding, and marketing, is a vital research area with substantial industry impact.
-As discussed in [13, 9, 12, 21, 8, 23, 4, 46], a well-crafted graphic design can effectively convey messages through a skillful blend of shapes, colors, and fonts, all while respecting design principles and aesthetics.
-A significant amount of research has been invested in generating specific components of professional graphic design images, such as layouts or typography attributes.
-For instance, FlexDM [16, 43] leverages a unified Transformer architecture for element prediction.
-Additional research [30] focuses on typography generation, while UI-Diffuser [35] investigates mobile UI generation.
-On the other hand, the majority of studies [12, 23, 4, 46] primarily concentrate on layout generation, assuming that all other information is at hand.
-Unlike the above-mentioned methods that focus on individual aspects within the graphic design generation task, this study aims to develop an autonomous, comprehensive system capable of executing all design generation tasks from clear user intent descriptions.
+Autonomous Agent based on LLMs Recent studies [44, 19, 38, 43, 28, 11, 27, 39, 7, 39, 21] have proposed the development of advanced (multi-)agent systems leveraging powerful LLMs to tackle various tasks.
+Some of them [44, 19, 38, 43, 28] concentrate more on image generation tasks.
+Our model primarily distinguishes itself from these works in the mechanism of task decomposition and the usage of tools.
+While most existing studies incorporate tool usage knowledge by designing tailored prompt instructions or in-context examples for GPT-/, or GPT-4V(ision) to iteratively refine the text prompts, our Cole decomposes the complicated graphic design image generation task into a hierarchical set of generation tasks, and performs dedicated fine-tuning of each generation model on our constructed professional graphic design dataset.
+Design Image Generation The majority of studies [12, 22, 5, 45, 29, 34, 15, 42] primarily concentrate on individual aspects within the graphic design generation task, assuming that all other information is at hand.
+Unlike the these methods, this study aims to develop an autonomous, comprehensive system capable of executing all design generation tasks from clear user intent descriptions.
 This task formulation presents a significantly greater challenge compared to most existing approaches.
@@
 Each of these roles is designed to address specific aspects of the graphic design process, as indicated in the references [35, 33].
-Consequently, we have broken down the text-to-design generation task into a hierarchical structure of tasks.
-These consist of the intention-to-JSON generation task, text-to-background image generation task, text-to-object image and alpha mask generation task, typography attribute generation task, and the graphic rendering task.
-The tasks of intention-to-JSON generation and typography attribute generation both follow a hierarchical structure.
-They necessitate the prediction of missing values given a predefined structured request.
-We introduce the comprehensive hierarchical framework of our Cole system as depicted in Figure 3.
-To effectively train the specialized design-agent system, we initially amass approximately 100,000 high-quality raw graphic design images from the internet.
-Following this, we organize these graphic design data into a hierarchical structure which includes components such as (intention, JSON) pairs, (caption, background-image) pairs, (caption, object-image) pairs, and (text-image, typography JSON) pairs.
+Consequently, we train an LLM for multi-layered task planning, breaking down the text-to-design generation into a hierarchical structure of tasks.
+These include the intention-to-JSON generation, text-to-background image generation, text-to-object image and alpha mask generation, typography attribute generation, and the multi-layered graphic editing and rendering tasks.
+The tasks of intention-to-JSON and typography attribute generation both adhere to a hierarchical structure, requiring the prediction of missing values based on a predefined structured request.
+We introduce the comprehensive hierarchical framework of our Cole system, as depicted in Figure 4.
+To effectively train the Cole system, we initially amass approximately 100,000 high-quality raw graphic design images from the internet, covering multiple graphic design sources.
+Following this, we organize the graphic design data into a hierarchical structure that includes components such as (intention, JSON) pairs, (caption, background-image) pairs, (caption, object-image) pairs, and (text-image, typography JSON) pairs.
 We will further explore the process of constructing each component in the ensuing discussion.
-3.2 Design LLM
-We first formulate the intention-to-JSON prediction task as a masked field prediction problem following [15].
-This approach learns to predict element types, positions, attributes, and text using a unified Transformer architecture.
+3.2 Design LLM: Intention Recaption and Layout Planning
+We first detail the process of generating (intention, JSON) pair data for graphic design and then demonstrate the implementation of the Design LLM, aimed at intention recapitulation and planning multi-layered layouts.
+Creating Intention for Design Image Given that both intentions and detailed captions are absent in the raw graphic design data, we first employ GPT- to generate the intentions from the raw image information.
+This includes the image title, format, keywords, and all visual text present in the image.
+This is accomplished with the aid of the system prompt displayed in the “Intention Generation Prompt” (shown in Supplementary).
+The left of Figure 5 presents a representative example, showcasing the ground-truth pairs of graphic design images and the generated user intentions.
+We generate the intention for nearly graphic design images.
+Creating JSON for Design Image Subsequently, we prepare the structured JSON data to represent the design images, which include multiple key-value pairs such as ‘global caption’, ‘category’, ‘keywords’, and ‘layers’.
+Within ‘layers’, we further decompose it into three sub-layers: ‘background layer’, ‘object layer’, and ‘text layer’.
+Considering that not all design images contain the object layer, we also use a ‘flag’ to mark whether the object layer exist.
+For all captions, we opt to utilize the recently developed large multimodal model known as the LLaVA-1.5-13B model [24], which is effective in generating detailed captions for the entire image, including background and object image captions.
+The right side of Figure 5 presents a representative example, showcasing the generated JSON file for the given image.
+We create matching JSON files for nearly graphic design images, each paired with the above generated intentions.
+Implementation and Training Settings After constructing the intention and JSON pairs for the graphic design images, we formulate the intention-to-JSON prediction task as a masked field prediction problem following [14], which is very different from the conventional text-to-text prediction task.
+This approach learns to translate intentions into detailed JSON files and conducts multi-layered layout planning by predicting the global captions, background layer captions, object layer captions, contents of the text layers including the heading, sub-heading, and body-text.
 Furthermore, we adopt a strategy from [1] to construct causally masked multi-modal LLMs by leveraging the Llama-B [32].
 We have fine-tuned the causally masked Llama-B model using our assembled dataset of nearly (intention, JSON) pairs, over an approximate span of epochs.
-Our empirical findings suggest that incorporating additional predictions of the complete caption, along with key words from the graphic design image, facilitates subsequent tasks of background and object image generation.
-Given that both intentions and detailed captions are absent in the raw graphic design data, we first employ GPT- and GPT- models to generate the intentions.
-This is accomplished with the aid of the system prompt displayed in the “Intention Generation Prompt”.
-For a more comprehensive understanding, we have included detailed intention generation prompt with in-context examples in the supplementary materials.
-Subsequently, we utilize the recently developed, large multimodal model known as LLaVA--B model [24].
-This model is effective in generating detailed captions for the entire image, background image, and object image captions.
-Figure 4 presents a representative example, showcasing the ground-truth triplets of graphic design images, alongside the generated user intentions and corresponding JSON files.
-In addition, we have also investigated the potential of the highly proficient GPT- model for directly generating JSON from user intentions.
-This is achieved by formulating an instructive prompt.
-Our empirical analysis, detailed in the experiment section, demonstrates that our methodology consistently surpasses the performance of GPT-.
-3.3 Text-to-Background Diffusion Model
-Given the current unavailability of mature text-to-SVG technology and the scarcity of high-quality text-SVG pairs, which leads to the absence of strong text-to-SVG generation models [17, 47, 41, 42], our study suggests consolidating all SVG elements and additional embellishments into one unified image layer.
-Therefore, we first extract the generated detailed description of the background image layer from the JSON file.
-Then we train a text-to-background diffusion model to generate the coherent background images consisting of all the SVG files and decoration elements.
-We opt to fine-tune our collected less than (caption, background image) pairs using the recent DeepFloyd IF [19] diffusion models for epochs.
+Our empirical findings suggest that incorporating additional predictions of the ‘global caption’, along with ‘keywords’ from the graphic design image, facilitates subsequent tasks of background and object image generation.
+3.3 Text-to-Background Diffusion Model: Visual Planning to Generate Canvas Placeholders
+We now present the methodology for preparing (caption, background-image) pairs and delve into the complexities of training a dependable background image to serve as a canvas placeholder.
+This involves strategically planning about where to allocate sufficient space for the seamless integration of diverse elements.
+Preparing Text-Background Pairs Given the current unavailability of mature text-to-SVG technology and the scarcity of high-quality text-SVG pairs, which leads to the absence of strong text-to-SVG generation models [16, 46, 40, 41], our study suggests consolidating all SVG elements and additional embellishments into one unified background image layer.
+Then we extract the generated detailed description of the unified background image layer from the JSON file.
+The left side of Figure 6 displays an example of the decomposed background image layer from a given graphic design image.
+We construct less than (caption, background image) pairs for use as training data.
+Implementation and Training Settings Then we train a text-to-background diffusion model to generate the coherent background images consisting of all the SVG files and decoration elements.
+We opt to fine-tune our collected (caption, background image) pairs using the DeepFloyd IF [18] diffusion models for epochs.
 The output resolution is .
 Following this, we employ the generated background images as conditioning input for the text-to-object diffusion model, which predicts coherent high-resolution object image layers based on their corresponding object prompts.
-3.4 Text-to-Object Diffusion Model
-Subsequently, we propose training a text-to-object diffusion model capable of perceiving both the object captions present in the JSON file and the visual pixels predicted by the text-to-background diffusion models.
-Following the instruct-P2P approach [3], we concatenate the noise latent vectors with the latent vectors of the conditioning background image with the alpha mask for the object layer, and we add additional input channels to the first convolutional layer.
+One key challenge in generating the background image lies in ensuring it serves effectively as a canvas placeholder.
+This requires reserving adequate space for incorporating various elements, including object image layers and text layers.
+Consequently, the text-to-background generation model must perform visual planning to determine where to allocate space.
+This approach significantly diverges from conventional methods like DALLE3, which often generate objects occupying the entire image space.
+We provide a qualitative comparison of the background images generated by different methods in the appendix.
+Another observed challenge is the length of the generated detailed captions, which can extend to tens of words, resulting in hundreds of text tokens - far exceeding the original default maximum length.
+To overcome this, we have increased the maximum number of text tokens from to .
+Offset-noise is important for background image generation We observe the graphic design background images tend to have a uniform color distribution and this aspect sets them apart from the pre-trained distributions of photorealistic images.
+Our empirical findings emphasize the effectiveness of the offset noise technique [23, 9] for high-quality background image generation.
+This is because, even when , the model’s input isn’t purely noise.
+The leaked signal carries low-frequency information, such as the overall mean of each channel.
+The implementation details are as follows: .
+We choose by default.
+3.4 Text-to-Object Diffusion Model: Visual Reasoning based on the Generated Background Image
+This section outlines the method for creating (caption, object image) pair data and showcases the deployment of the text-to-object diffusion model.
+This model is designed to produce an object layer that is not only stylistically coherent but also logically positioned.
+A key technical insight in this endeavor is the model’s need to predict an additional composite image, a step that is essential for ensuring the precision of the object placements.
+Preparing Text-Object-Background-AlphaMask Quadruples To train a text-to-object diffusion model capable of perceiving both the object captions present in the JSON file and the visual pixels predicted by the text-to-background diffusion models, we prepare the (object-caption, background-image, object-image, alpha-mask) quadruples.
+We extract the object-image and the alpha-mask from the raw data, while the object-caption and the background-image are prepared following the above-mentioned scheme.
+Implementation and Training Settings Following the instruct-P2P approach [3], we concatenate the noise latent vectors with the latent vectors of the conditioning background image with the alpha mask for the object layer, and we add additional input channels to the first convolutional layer.
 All available weights of the diffusion model are initialized from pre-trained checkpoints, while weights operating on the newly added input channels are initialized to zero.
@@
 Consequently, we can seamlessly blend the object image layer with the background image layer in accordance with the alpha mask.
-Empirically, we found that predicting the additional composed image is pivotal to ensuring a more reasonable placement of the object pixels.
-Therefore, our text-to-object diffusion model needs to predict channels.
-For more details, please refer to the experimental section.
-Similar to the text-to-background generation model, we optimize our assembled dataset, comprising approximately caption and object image pairs.
-The relatively smaller number of pairs is due to the frequent absence of the object image layer in many graphic design images.
+To ensure a more reasonable placement of object pixels and coherence in style with the background image, we propose predicting an additional composed image.
+This is achieved by adding three extra channels to both the input noise vectors and the output prediction, a step that proves critical for accurate object layer prediction.
+Therefore, our text-to-object diffusion model needs to predict channels, where the object image consists of channels, the object mask consists of channel, and the composed image consists of channels.
+Besides, we also increased the maximum number of text tokens from to following the text-to-background diffusion model.
+Similar to the text-to-background generation model, we optimize our assembled dataset, comprising approximately (object-caption, background-image, object-image, alpha-mask) quadruples.
+The relatively smaller number of quadruples is due to the frequent absence of the object image layer in many graphic design images.
 We employ the instruct-P2P scheme for optimization, which is based on the DeepFloyd [18] IF diffusion models, and this process is carried out over an estimated span of epochs.
-Figure 5 provides a selection of representative examples that demonstrate our method of decomposing complex graphic design images into distinct layers of background and object images.
+Qualitative Examples Figure 6 provides a selection of representative examples that demonstrate our method of decomposing complex graphic design images into distinct layers of background and object images.
 It’s important to note that predicting the object alpha mask is essential for facilitating flexible editing operations on the object pixels.
-However, ensuring accurate alpha mask prediction presents a complex challenge.
-Please refer to the supplementary materials for more in-depth technical details.
-3.5 Typography LMM
+3.5 Typography LMM: Layout Planning and Attribute Reasoning for Visual Text
 Typography [37] is the art and technique of arranging type to make written language legible, clear, and visually engaging for the reader.
-More succinctly, typography refers to the style and appearance of text.
+It therefore requires complex reasoning and planning capabilities.
 It encompasses the design of font style, appearance, and structure, with the goal of eliciting specific emotions and conveying certain messages.
 In essence, typography brings text to life.
-Contrary to recent efforts, such as Tanveer et al.
-[32], Iluz et al.
-[14], and He et al.
-[10], which primarily employ diffusion models to generate novel letter-level or word-level semantic or artistic typographies, our focus is more directed towards the utilization of an LMM (Large Multimodal Model).
+Contrary to recent efforts, such as [31], [13], and [10], which primarily employ diffusion models to generate novel letter-level or word-level semantic or artistic typographies, our focus is more directed towards the utilization of an LMM (Large Multimodal Model).
 This approach allows us to address the highly challenging task of image-level typography.
-This task involves the selection of font types, establishment of hierarchy arrangements, adjustment of leadings (line spacing), tracking (character spacing), and kernings (space between specific characters).
-In our experiments, we elected to fine-tune the LLaVA--B model [24] using nearly triplets for epochs.
-These triplets consist of multiple texts, composed images of backgrounds and objects, and typography information.
-Figure 6 displays a representative example of a Typography JSON file.
+This task encompasses selecting font types, establishing hierarchical arrangements, and adjusting leading (line spacing), tracking (character spacing), and kerning (space between specific characters) based on the composed images.
+Consequently, it demands a robust capability for layout planning and attribute reasoning from the Typography LMM.
+Preparing Composed Images and Typography Data We construct nearly triplets of data, each consists of multiple texts, composed images of backgrounds and objects, and typography information.
+Figure 7 displays a representative example of a typography JSON file, illustrating that it comprises attributes, with the value of ‘text’ already predicted by the Design-LLM.
+We extract all these typography attributes directly from the raw data of each graphic design image.
+Implementation and Training Settings In our experiments, we elected to fine-tune the LLaVA--B model [24] using nearly triplets for epochs.
 It is evident that predicting visually appealing typography information for each text box, based on the input image, is a complex task.
-This is due to the requirement of predicting approximately attributes for each box.
+This is due to the requirement of predicting approximately attributes for each text box.
 To our knowledge, even the most recent effort [29] only manages to predict typography attributes, assuming the other attributes are fixed.
@@
 We explain more details of the bin partition scheme in the supplementary material.
-3.6 Graphic Render
-As demonstrated in the preceding sections, we have successfully predicted the JSON file, complete with detailed captions and text elements.
-We have also generated the background and object image layers (with object alpha mask), ensuring they adhere to the principles of design aesthetics and layout.
-In addition, we have produced the typography JSON file, which precisely outlines the style and appearance of the text elements.
-We implement a graphic renderer founded on Skia [31], an open-source 2D graphics library.
-It is important to highlight that the rendering quality, as per the official implementation of CanvasVAE [43], presents significant potential for enhancement.
-Figure 7 showcases a qualitative comparison of the results obtained from our rendering implementation, contrasted with those generated by CanvasVAE.
-Upon evaluation, it is evident that our graphic renderer significantly outperforms the open-sourced CanvasVAE.
-Original
-Ours
-CanvasVAE
-Original
-Ours
-CanvasVAE
+3.6 Multi-Layered SVG Editor and Renderer: Support Layer-wise Flexible User Editing
+After generating all the design image layers, we propose compiling them into a single SVG file using the open-source tool SVG-Edit[30].
+Following this, we opt to utilize Justinmind222https://www.justinmind.com/, a tool that facilitates a wide range of flexible, layer-wise user editing operations.
+These operations include repositioning and resizing text boxes or object layers.
+Additionally, Justinmind allows for modifications to be made to font size, color, type, and other attributes within each text box, enhancing the user’s ability to customize their designs further.
+We present the individual image layers and the fully composed image in Figure 1.
+Additionally, we showcase a series of representative examples that demonstrate the outcomes following various layer-wise editing operations in Figure 9.
 3.7 Reflect LMM & Quality LMM
-We find that initial graphic design outputs may not always align with user expectations.
-Therefore, we propose the integration of an added Reflect-LMM.
+Considering that initial multi-layered graphic design outputs may not always align with user expectations, we further propose the integration of an added Reflect-LMM.
 This component is designed to critically analyze the generated design images and predict necessary adjustments to the JSON file, thereby further improving the resulting graphic designs.
-The entire self-reflection framework is illustrated on the left side of Figure 8.
-As an integral part of our implementation plan, we fine-tune a LLaVA--B model on a dataset that comprises more than artificially generated noisy JSON files.
-Each of these files is paired with a corresponding ground-truth JSON file.
-Our primary modification involves the addition of random shift noise to the position of each ground-truth typography block.
-Besides, we utilize GPT-4V(ision) [25] to evaluate the quality of the generated graphic design images, as displayed in the “Quality Assurance Prompt for GPT-4V(ision)” in the appendix.
-Through empirical analysis, we have found that GPT-4V(ision) excels at evaluating the quality of main aspects such as design and layout, content relevance, graphics and images, and innovation.
-However, it falls short in assessing the quality of typography and color.
-In addition, we also incorporate the evaluation prompt design as introduced in a recent concurrent effort [18].
-This evaluates the graphic design images from approximately aspects, including text rendering quality, composition and layout, color harmony, cinematography, style, image-text alignment, aesthetics, and overall design.
-4 Experiment
-(a) of text box, words, chars
-(b) Wordcloud
 | Category |
@@
 |
-(a) covers & headers
+Reflect LMM The entire self-reflection framework is illustrated on the left side of Figure 9.
+As an integral part of our implementation plan, we fine-tune a LLaVA--B model on a dataset that comprises more than artificially generated noisy JSON files.
+Each of these files is paired with a corresponding ground-truth JSON file.
+Our primary modification involves the addition of random shift noise to the position of each ground-truth typography block.
+The key insight lies in modeling the conditional transformation of typography attributes from suboptimal to ideal, leveraging the understanding of the rendered composite images.
+Quality LMM We utilize GPT-4V(ision) [25] to evaluate the quality of the generated graphic design images, as displayed in the “Quality Assurance Prompt for GPT-4V(ision)” in the appendix.
+Through empirical analysis, we have found that GPT-4V(ision) excels at evaluating the quality of main aspects such as design and layout, content relevance, graphics and images, and innovation.
+However, it falls short in assessing the quality of typography and color.
+In addition, we also incorporate the evaluation prompt design as introduced in a recent concurrent effort [17].
+This evaluates the graphic design images from approximately aspects, including text rendering quality, composition and layout, color harmony, cinematography, style, image-text alignment, aesthetics, and overall design.
+4 Experiment
+(a) covers & head.
 (b) marketing
@@
 (e) advertising
-background image
-object image
-object alpha mask
-composed image
-To assess our Cole’s proficiency in generating graphic design images, we first build a DesignerIntention benchmark comprising over approximately graphic design intention prompts.
-These prompts serve as a tool to evaluate the model’s performance from diverse perspectives.
-Second, as mentioned in the previous section, we propose utilizing GPT-4V(ision) to evaluate the quality of the generated graphic design images and compare our system with the latest systems including DeepFloyd/IF(+GPT-), SDXL(+GPT-), and DALLE3 (+GPT-) following [2, 18].
+Ours
+Ours+Edit
+CanvaGPT
+DeepFloyd/IF
+SDXL
+DALLE3
+Seed1
+Seed2
+Seed3
+Seed4
+To assess our Cole’s capability for design-oriented planning and reasoning in generating graphic design images, we conducted extensive experiments on our proposed DesignerIntention, which consists of approximately graphic design intention prompts.
+Leveraging GPT-4V(ision), we conducted a multi-faceted comparison [2, 17] between our approach and the state-of-the-art image generators DeepFloyd/IF, SDXL, DALLE3, and CanvaGPT, which utilize GPT- as an intention interpreter.
+Subsequently, we assembled approximately 20 general users and 6 professional designers to conduct a subjective assessment using COLE, comparing it separately with both DALLE3 and CanvaGPT.
+Then we show that the zero-shot performance with our Typography-LMM even outperforms the latest layout planning methods in Crello text box placement task.
 Last, we carry out ablation experiments to scrutinize the performance of each component and discuss the limitations of our system.
@@
 In a complementary effort, we have compiled an extra set of around unique, creative user intentions.
-The graphic design images derived from these inventive prompts are displayed for your reference in last row of Figure 16.
-Refer to the supplementary for an all-inclusive review of these roughly intention prompts, along with the creative ones.
-Figure 9 furnishes comprehensive statistics of our benchmark from myriad perspectives.
+Refer to the supplementary material for an all-inclusive review of these approximately intention prompts, along with the approximately creative ones, and comprehensive statistics.
 We conduct all the experiments on DesignerIntention by default.
 4.2 Main Results
-Comparison with State-of-the-art
-Figure 15 presents a comparison with various state-of-the-art text-to-image generation models that are augmented with GPT-.
-The usage of GPT- stems from the fact that traditional text-to-image models generally necessitate detailed descriptions, which are often missing in the intention prompt.
+Comparison with State-of-the-art Figure 11 presents a comparison with various state-of-the-art text-to-image generation models that are augmented with GPT-.
+Notably, we also include the results of the widely recognized custom CanvaGPT [4], which integrates GPT-4 with Canva’s templates.
+Each Canva template fully produced by professional designers incurs significant costs.
+The usage of GPT- stems from the fact that both traditional text-to-image models and CanvaGPT generally necessitate detailed descriptions, which are often missing in the intention prompt.
 Consequently, we utilize GPT- across all models to translate the user’s simple intentions into intricate captions.
@@
 In addition, we showcase their predicted numerical quality scores from GPT-4V(ision) based on quality assurance prompt in Figure 11.
-A careful evaluation reveals that our Cole achieves very competitive quality accessing results even compared to the latest DALLE3.
-Our Cole achieves the best quality when generating covers & headers or posters.
+A careful evaluation reveals that our Cole achieves very competitive quality accessing results even compared to the latest DALLE3 and CanvaGPT.
+Our Cole secures leading performance across three critical dimensions: design layout, typography color, and innovation, consistently achieving second-best in quality.
 An interesting observation is that the background image generated by DALLE3 is of exceptional quality but the text regions are non-editable.
-We also empirically show some impressive results of our Cole when replacing our text-to-background generation model with DALLE3 by leveraging the text segmentation models, inpainting models, and others.
-Seed #1
-Seed #2
-Seed #3
-Seed #4
-Seed #5
-initial image
-+ object editing
-+ text editing
-+ combined editing
-Ours
-DeepFloyd/IF
-SDXL
-DALLE3
-Qualitative results
-Figure 16 showcases a variety of graphic design images entirely created by our system.
+Figure 1 shows some impressive results of our Cole when replacing our text-to-background generation model with DALLE3 by leveraging the text segmentation models, inpainting models, and others.
+Qualitative results Figure 12 showcases a variety of graphic design images entirely created by our Cole system.
 It’s noteworthy that our system performs exceptionally well, considering that each model within it is trained independently in a hierarchically organized manner.
@@
 4.3 Ablation Experiments
-Design LLM
-We assess the performance of the Design LLM by calculating the semantic textual similarity (STS) between the predicted text and the ground-truth ones using Sentence-BERT [28].
-This evaluation is done across various keys, such as global or background or object caption, keywords, and heading&sub-heading&bodytexts, following the generated JSON file.
-Their average STS scores are , , , , and , respectively.
-Text-to-Background & Text-to-Object
-Figure 12 showcases the generated background image layer, object image layer, and the composed image layer.
-The displayed visualizations highlight the proficiency of our text-to-object generation model in crafting a coherent object image layer and its corresponding alpha mask.
-We have also carried out a detailed comparison of the background images generated by our model and others [26, 19, 2].
-Refer to the supplementary material for more details.
-Typography LMM & Reflect LMM
-We visualize the predicted results with the ground-truth images as the condition in Figure 12.
+Text-to-Background & Text-to-Object Figure 15 illustrates the critical role of incorporating composed image prediction into the text-to-object diffusion model.
+This integration markedly enhances the model’s ability to plan layouts.
+Furthermore, we have conducted an exhaustive comparison between the background images produced by our model and those generated by other methods, such as those described in [26, 18, 2].
+We found that most alternative approaches often result in objects that fill the entire rectangular canvas.
+For additional information, please consult the supplementary material.
+Typography LMM & Reflect LMM We visualize the predicted results with the ground-truth images as the condition in Figure 14.
 We assess these results from two perspectives: typography scoring with GPT-4V(ision) and the localization accuracy including , , and .
-For a more comprehensive comparison and the ablation of the reflect-LMM, refer to the supplementary material.
-Design Diversity
-Designers often face the critical challenge of delivering diverse designs while maintaining high quality.
-Our Cole system empirically excels in this regard, as evidenced by its operation with various random seeds shown in Figure 14.
-Flexible Editing
-Our Cole offers a versatile and interactive platform that allows users to edit both the object image layer and typography.
-This functionality augments the capability to manipulate generated graphic design images from various perspectives.
-Figure 14 showcases examples of different editing operations.
-For instance, users have the flexibility to resize, reposition, or replace the object image layer, as well as to freely modify the font type and size.
+To assess the generalization capability of our approach in layout planning and attribute reasoning for visual text, we evaluate the zero-shot performance on the Crello benchmark and summarize the comparison results in Table 3.
+Our Typography-LMM outperforms the previous model, FlexDM, by +4.5 and +6.2 in terms of IoU score for single and multiple text box placement tasks, respectively Refer to the supplementary material for the ablation of reflect-LMM.
+Design Diversity Designers frequently confront the challenge of producing varied yet high-quality designs.
+Our Cole system demonstrates exceptional performance in this aspect, as illustrated in Figure 14.
+We provide additional ablation experiments on the Design LLM and others in the supplementary material.
 5 Conclusion
-This paper introduces an exceptionally efficient hierarchical generation framework., Cole, that simplifies the intricate process of graphic design generation by breaking it down into manageable, coordinated sub-tasks, each yielding distinct outputs.
-Leveraging the benefits of this hierarchical task decomposition, our system can generate competitive graphic design images from simple user intent even when compared to DALLE3 on the challenging DesignerIntention benchmark.
-The core strengths of our Cole system lie in ensuring precise text rendering and providing flexible editing options.
-We have also pinpointed several key limitations of our Cole system in the context of generating graphic design images, and we plan to tackle these in future research.
+This paper introduces an exceptionally efficient hierarchical generation framework, Cole, that simplifies the intricate process of graphic design generation by breaking it down into manageable, coordinated sub-tasks, each yielding distinct outputs.
+By leveraging design-oriented visual planning, reasoning, and layer-wise generation capabilities during this hierarchical task decomposition, our system can generate multi-layered and editable graphic design images.
+We demonstrate the advantages of our approach by comparing it to the most powerful systems like DALLE3 and CanvaGPT.
+Additionally, we have pinpointed several key limitations of our Cole system in the context of generating graphic design images, and we plan to address these in future research.
 Acknowledgement We extend our gratitude to Xinyang Li for his invaluable assistance with data preparation.
-Our thanks also go to Yang Ou for providing insightful comments from a professional designer’s perspective.
+Additionally, we are grateful to Yang Ou, along with both designers and non-designers, for their insightful feedback from a professional design and normal user perspective.
 We appreciate the enlightening discussions offered by Chunyu Wang, Jianmin Bao, Mingxi Cheng, and Chong Luo.
@@
 Supplementary Material
-A Graphic Design Pipeline
-The generation of high-quality graphic designs from user intentions requires the harmonious integration of several roles or components, each tailored to address specific aspects of the graphic design process [36, 34].
-We initially offer a detailed plan that outlines these roles, clarifies their purposes, and explains their interactions.
-Figure 17 illustrates the complete pipeline.
-Creative Design and Interpretation
-The primary function of this role involves understanding user intentions and translating them into tangible design elements like titles, subtitles, headings, subheadings, and body text.
-It also sets the tone and style of the graphic design, which incorporates elements such as primary color, theme, style, and captions.
-Visual Composition and Enhancement
-The key responsibility of this role is to uphold the principles of design aesthetics and layout, which involves the generation, assessment, and adjustment of visual appeal and spatial organization.
+A Graphic Design Pipeline of Human Designer
+The creation of high-quality graphic designs from user intentions necessitates a seamless integration of diverse roles or components, each uniquely designed to tackle specific facets of the graphic design process [35, 33].
+Initially, we present an elaborate blueprint that delineates these roles, elucidates their functions, and delineates their interactions.
+Figure 16 depicts this comprehensive pipeline, serving as the conceptual foundation for the design of our Cole system.
+Creative Design and Interpretation The core responsibility of this role is to interpret user intentions and convert them into specific design elements, including titles, subtitles, headings, subheadings, and body text.
+This role also establishes the tone and style of the graphic design, encompassing aspects like primary color, theme, style, and captions.
+Visual Composition and Enhancement The key responsibility of this role is to uphold the principles of design aesthetics and layout, which involves the generation, assessment, and adjustment of visual appeal and spatial organization.
 The generation process mainly utilizes a comprehensive library of images, icons, and textures to enhance the design.
@@
 For instance, it is imperative to not only generate images of objects, decorations, and backgrounds, but also to predict various typographic attributes for elements such as the heading, subheading, and body text.
-Feedback and Quality Assurance System
-The principal function of this role is to integrate user feedback into the design in an iterative manner, with the objective of ensuring that generated designs comply with quality benchmarks and regulatory standards.
+Feedback and Quality Assurance System The principal function of this role is to integrate user feedback into the design in an iterative manner, with the objective of ensuring that generated designs comply with quality benchmarks and regulatory standards.
 This module guarantees that the designs not only align with user expectations, but also fulfill professional and legal criteria.
 It maintains a continuous interaction with the first two modules to guide the design refinement process.
-Collaboration Loop
-This system ensures that the designs not only meet user expectations, but also adhere to professional and legal standards.
+Collaboration Loop This system ensures that the designs not only meet user expectations, but also adhere to professional and legal standards.
 It maintains a constant dialogue with the first two modules to navigate the design refinement process effectively.
-Discussion
-Unlike traditional graphic design systems that rely significantly on the expertise of professional designers, we aspire to create a DesignerAgent.
+Discussion Unlike traditional graphic design systems that rely significantly on the expertise of professional designers, we aspire to create a DesignerAgent.
 This innovative system is designed to generate superior graphic designs, leveraging user intention descriptions alone.
 By fine-tuning state-of-the-art Large Multimodal Models (LMMs), Large Language Models (LLMs), and diffusion models, we aim to fully exploit our carefully curated, high-quality graphic design datasets to achieve this objective.
-B. Detailed System Prompts for GPT-3.5 and GPT-4V(ision)
-We introduce the Intention Generation Prompt for GPT3.5, Quality Assurance Prompt for GPT-4V(ision), and Pair-wise Comparison Prompt for GPT-4V(ision) as follows: The Intention Generation Prompt is crafted to deduce the user’s purpose utilizing diverse image data, such as keywords, all text visible in the image, the image’s title, among others.
+B Detailed System Prompts for GPT-3.5 and GPT-4V(ision)
+We introduce the Intention Generation Prompt for GPT-3.5, Quality Assurance Prompt for GPT-4V(ision), and Pair-wise Comparison Prompt for GPT-4V(ision) as follows: The Intention Generation Prompt is crafted to deduce the user’s purpose utilizing diverse image data, such as keywords, all text visible in the image, the image’s title, among others.
 The Quality Assurance Prompt for GPT-4V(ision) is established to evaluate the quality of the created graphic design images across roughly five dimensions.
@@
 The Pair-wise Comparison Prompt for GPT-4V(ision) is proposed to objectively compare our generated images with results from other methods from six perspectives of graphic design.
-C. The Detailed List of DesignerIntention
-A comprehensive list of designer intention prompts is summarized across Table 4, Table 4, Table 4, and Table 6.
-Table 4 details the intention prompts corresponding to the images exhibited in Figure 1 of the main paper.
-In Figure 20, we show a few example pairs that illustrate the original image information alongside the corresponding generated intention prompt.
-Table 4 presents the intention prompts augmented with GPT-, which were utilized for DALLE3 in Figure 2 of the main paper.
-In Table 4, we display the augmented intention prompts with GPT- that were employed for DeepFloyd/IF, SDXL, and DALLE3 in Figure 10 of the main paper.
-Last, Table 6 provides the detailed intention prompts for the design images generated, as depicted in Figure 11 of the main paper.
-D. More Ablation Experiments
+C The Detailed Prompt List of DesignerIntention
+A comprehensive list of designer intention prompts is summarized across Table 6, Table 6, Table 7 and Table 8.
+Table 6 details the GPT-4 augmented intention prompts corresponding to the images exhibited in Figure 1 of the main paper.
+Table 6 presents the intention prompts which were utilized for Cole in Figure 2 of the main paper.
+Table 7 presents the intention prompts and GPT-4 augmented intension prompts which were utilized for Cole in Figure 11 of the main paper.
+Last, Table 8 presents the intention prompts which were utilized for Cole in Figure 12 of the main paper.
+D More Details of Bin Partition Scheme within Typography-LMM
+To streamline the Typography-LMM’s prediction process, we adopt a discretization strategy, i.e., bin partition scheme, for continuous floating-point values based on the approach described in [6].
+This involves categorizing continuous variables into discrete bins and representing the values within each bin with a unique token.
+The specifics of our bin partitioning scheme are detailed below:
+Font Size: The range [2, 200] is divided into 100 sub-intervals.
+Angle: The range [0, 2] is segmented into 64 sub-intervals.
+Color (RGB values): The range [0, 255] is split into 32 sub-intervals.
+Text Bounding Box Coordinates: The normalized coordinates within the range [-1, 1] are partitioned into 256 subspaces.
+Opacity: The range [0, 255] is divided into 8 subspaces.
+Letter/Line Spacing: The normalized values within the range [0, 1] are divided into 40 sub-intervals, adjusted according to the image width.
+We estimate the start and end points of all the above partitions based on the statistics derived from the training set, accommodating a broad spectrum of typography characteristics, including font size, angle, letter and line spacing, opacity, dimensions of the font bounding box (left, top, width, height), and color.
+E More Ablation Experiments
+Design LLM We calculate the semantic scores (semantic textual similarity scores estimated based on Sentence-BERT) to compare the ground-truth with predictions from our Design-LLM, GPT-3.5, and GPT-4 across the test dataset.
+To ensure accurate predictions, we supply three in-context examples for both GPT-3.5 and GPT-4.
+Furthermore, we restrict GPT-3.5 and GPT-4 to generate values for only one key query at a time, simplifying the task’s complexity.
+The results, detailed in Table 4, reveal that our Design-LLM significantly outperforms both GPT-3.5 and GPT-4.
+| Keys in JSON | Design LLM | GPT-3.5 | GPT-4 |
+| Global Caption | 68.57 | 56.31 | 59.56 |
+| Category | 80.32 | 80.93 | 78.65 |
+| Keywords | 89.31 | 84.67 | 84.14 |
+| Background Caption | 57.69 | 48.06 | 52.37 |
+| Object Caption | 73.59 | 73.21 | 73.16 |
+| Text layer (heading, sub-heading, body-text) | 54.75 | 46.80 | 53.56 |
 Text-to-Background & Text-to-Object To assess the efficacy of our text-to-background diffusion model, we conducted a comparative analysis involving DeepFolyd/IF, SDXL, and DALLE3, as depicted in Figure 20.
@@
 Our findings reveal a notable trend among current state-of-the-art diffusion models: they often generate objects that occupy the entirety of the image, neglecting adequate space for accommodating visual text.
-| Category |
-Intention Prompt |
-| creative | Design an animated digital poster for a fantasy concert where Joe Hisaishi performs on a grand piano on a moonlit beach.
-The scene should be magical, with bioluminescent waves and a starry sky.
-The poster should invite viewers to an exclusive virtual event.
-|
-| advertising | Create a customer testimonial advertisement showcasing a pink clothing collection. The advertisement includes a positive review from Jenny Wilson. |
-| marketing | Create a poster advertising a corner spice shop with a wide range of spices and ingredients. |
-| posts | Promote a stand-up comedy show happening on November 12th, 20xx. |
-| marketing | Design a gift certificate for a travel insurance company promoting their services with a 30% discount offer.
-The certificate should convey a sense of safety and protection while travelling.
-|
-| posts | Design a recruiting post for a graphic designer job opening. Encourage interested candidates to send their CV to yourname@email.com |
-| marketing | Design a Christmas sale poster with a fun and festive look. The poster should advertise up to 50% off on Christmas accessories, with a limited time offer. |
-| marketing | Design a gift certificate for a travel insurance company that offers a 30% discount for all kinds of travel. |
-| creative | Design an alluring poster for a blues concert in a clearing of the Amazon Rainforest.
-The artwork should evoke the soulful tunes of blues music echoing through the wilderness.
-Use a moody, midnight blue palette and incorporate silhouettes of blues instruments and nocturnal wildlife.
-|
-| posts | Create a LinkedIn post with a cartoon illustration to promote an article about the benefits and growing popularity of remote work from home.
-The post links to a website for more information about the topic and is associated with a specific company and CEO John Doe.
-|
-| marketing | Create a poster advertising babysitting services.
-The design should be warm and inviting, featuring a little girl and emphasizing professional and reliable childcare.
-The business name and phone number should be prominently displayed (+123 456 7890).
-|
-| marketing | Create a poster advertisement for a restaurant’s happy hour special, showcasing the luxury and elegance of the establishment.
-The poster should include the restaurant’s name, location, website, and phone number for reservations and delivery.
-Happy Hour runs from 7 PM to 11 PM.
-|
-| marketing | Create a real estate rental property flyer with a cozy and inviting interior design.
-Showcase the apartment’s features, amenities, and price.
-Include contact information for the real estate agency and their website address.
-The monthly rental price is $13,950.
-Bedrooms: 2, Bathrooms: 3, Square Footage: 1500.
-|
-| creative | Design an interactive digital invitation for a space-themed gala.
-The invitation should feature an animated rocket launch sequence and include a scannable QR code that plays a personalized greeting from a holographic astronaut.
-|
-| creative | Create a series of posters for a fictional Ancient Greek Olympics, with each poster highlighting a different sport like chariot racing or discus throw, styled in traditional Greek pottery art.
-|
-| covers& headers | Create an Instagram story promoting a special Christmas offer for a chocolate drink. The offer includes a 20% discount and encourages customers to order now. |
 | Category |
@@
 |
+| marketing | An elegantly designed gift certificate for a reputable travel insurance company.
+The certificate, made of high-quality paper, is adorned with motifs of different modes of travel - airplanes, ships, and trains.
+A shiny, golden 30% off stamp is prominently placed at the center.
+The company’s logo is tastefully positioned at the top, while the certificate’s edges are softly curved, enhancing its appeal.
+|
+| covers& headers | A vividly illuminated stage under a starry night sky, acting as the primary backdrop.
+Smack in the center is a caricature of John Oliver, his signature glasses glinting, holding a microphone with a wide grin.
+At the top, bold letters spell ”Laugh Out Loud with John Oliver”.
+Below, in elegant script, is the location ”At The Comedy Club, New York”.
+A chuckling audience silhouette completes the lively atmosphere.
+|
+| creative | A collection of vibrant posters for a fictitious Ancient Greek Olympics.
+Each poster features a different sport: a fiercely competitive chariot race with horses kicking up dust, a discus thrower showcasing his strength against a setting sun, all stylized like traditional Greek pottery art, with the sportsmen clad in traditional tunics, their muscles rippling under the warm, orange glow of the Grecian sun.
+The borders of each poster are adorned with intricate patterns of laurel wreaths.
+|
+| advertising | An eye-catching weekend sale advertisement for skincare products, featuring prominently a sleek bottle of cosmetic serum bathed in a soft, ethereal glow.
+A bold, ruby-red text proudly declares ”All Items Discounted Up To 30%”.
+Surrounding the serum are various skincare items, all discounted, framed against a backdrop of vibrant autumn leaves, reflecting this season’s promotion.
+|
 | Category |
-GPT4 Refined Intention Prompt |
-| advertising | A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged.
+Intention Prompt |
+| creative | Design an animated digital poster for a fantasy concert where Joe Hisaishi performs on a grand piano on a moonlit beach.
+The scene should be magical, with bioluminescent waves and a starry sky.
+The poster should invite viewers to an exclusive virtual event.
+|
+| creative | Design an alluring poster for a blues concert in a clearing of the Amazon Rainforest.
+The artwork should evoke the soulful tunes of blues music echoing through the wilderness.
+Use a moody, midnight blue palette and incorporate silhouettes of blues instruments and nocturnal wildlife.
+|
+| covers& headers | Create an Instagram story promoting a special Christmas offer for a chocolate drink. The offer includes a 20% discount and encourages customers to order now. |
+| advertising | Create a customer testimonial advertisement showcasing a pink clothing collection. The advertisement includes a positive review from Jenny Wilson. |
+| Category |
+Intension Prompt |
+GPT-4 Augmented Prompt |
+| advertising | Create a customer testimonial advertisement showcasing a pink clothing collection.
+The advertisement includes a positive review from Jenny Wilson.
+| A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged.
 In the corner, a gleaming starburst accentuates a glowing testimonial from Jenny Wilson, her words of praise shimmering.
@@
 |
-| events | A vibrant birthday card bathed in festive hues of red, orange, and yellow.
-A delectable cake, crowned with an assortment of juicy fruits, serves as the focal point.
-Around it, confetti and streamers tumble playfully.
-The card invites you to a birthday celebration on June 22nd, from 8pm to 11pm, in a quaint house located at 123 Street, Any City 4567.
+| events | Create a birthday card with a festive and colorful design featuring a cake with fruits.
+The card is an invitation to a birthday party happening on June 22nd from 8pm to 11pm at 123 st.
+Any city 4567.
+| A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged.
+In the corner, a gleaming starburst accentuates a glowing testimonial from Jenny Wilson, her words of praise shimmering.
+Beneath the testimonial, elegantly scripted contact information is highlighted: phone number 123-456-789, website WWW.FASHIONSTYLE.COM, and an address at 123 Anywhere ST., Any City.
 |
-| marketing | Picture a vibrant gift certificate for a leading travel insurance company.
-A globe subtly watermarked in the background, planes, and ships creatively darting across.
-Emblazoned on the center in bold, gold lettering is the irresistible offer: ”30% Discount on All Travel Insurance Policies”.
-Borders of the certificate are adorned with famous landmarks, adding a touch of wanderlust.
-The company’s logo sits proudly at the top, promising security and peace of mind for all journeys.
-|
-| posts | A vibrant autumnal landscape with leaves in shades of gold, amber, and crimson cascading from the trees.
-A sleek banner sweeps across the scene, displaying the text ”30% off Autumn Collection”.
-A stylish arrow gently points downwards towards a glowing ”Link in Bio” button.
-The whole image exudes an irresistible allure, beckoning Instagram viewers to partake in this exclusive offer.
-|
-| covers& headers | A vintage-inspired thumbnail for a vlogger episode, prominently featuring a curated assortment of five rare antique telephones, each unique in design and era.
-The collection is artfully arranged on a mahogany table against a backdrop of antiquated maps and time-worn books.
-The eye-catching title, ’Top 5 Antique Phones’, is elegantly superimposed on the scene in a classic, bold typeface.
-|
-| creative | Imagine a captivating poster for a blues concert nestled in a moonlit clearing of the Amazon Rainforest.
-The midnight blue canvas paints an enchanting scene of silhouetted blues instruments - a saxophone, a guitar, and a harmonica - intertwined with the silhouettes of nocturnal wildlife like sleek panthers and hooting owls.
-The artwork pulsates with the soulful rhythm of blues music, echoing through the wilderness, stirring the soul.
+| creative | Design an alluring poster for a blues concert in a clearing of the Amazon Rainforest.
+The artwork should evoke the soulful tunes of blues music echoing through the wilderness.
+Use a moody, midnight blue palette and incorporate silhouettes of blues instruments and nocturnal wildlife.
+| An enticing poster for a blues concert set in an Amazon Rainforest clearing.
+The artwork is bathed in a mysterious, midnight blue hue, resonating the soulful melody of blues music.
+The scene is punctuated by silhouettes of blues instruments - a haunting saxophone, a melancholic guitar, and a poignant piano.
+The wilderness comes alive in the shadows of nocturnal creatures - a stealthy jaguar, an elusive tree frog, and a silent owl perched on a branch, all harmoniously intertwined with the rhythm of the blues.
 |
@@
 Intention Prompt |
-| covers& headers | Create a Youtube thumbnail for a video showcasing how to make Italian dishes at home, specifically focusing on rolling out dough. |
-| covers& headers | Create a Youtube thumbnail to promote a pastry making tutorial series.
-The series is called ’The Best Pastry in the World’ and is hosted on the channel ’Your Channel’.
-The thumbnail should showcase a collage of homemade pastries, specifically croissants.
-Part 01 should be emphasized.
+| posts | Design a funny and creative Halloween Facebook post with humorous messages and stickers. Promote a Halloween party at Under Bar with the caption ’Get your treat, Trick or Treat, dude!’ |
+| marketing | Design a gift certificate for a travel insurance company that offers a 30% discount for all kinds of travel. |
+| covers& headers | Create a Youtube thumbnail for promoting an antique store subscription service. Showcase the variety of unique vintage goods available in the store. |
+| creative | Create a recruitment post for a company seeking an illustrator to join their team. |
+| marketing | Design a colorful store logo with an illustration of leaves to represent a fresh and vibrant shopping experience. The store may be named TB. |
+| posts | Create a Facebook post featuring a happy family with their child enjoying a leisurely day outdoors.
+Capture their stylish outfits and cute smile.
+Frame the image with a white border and add a shadow effect.
 |
-| covers& headers | Create a Youtube thumbnail for promoting an antique store subscription service. Showcase the variety of unique vintage goods available in the store. |
-| covers& headers | Design a Facebook event cover for a winter decoration workshop featuring lectures and demonstrations by Miranda Yates, Sam Knightly, Zachary Evans, and Veronica Floyd on December 24 at Lafayette Studio.
+| posts | Create a Facebook post to promote a special menu for pancakes at a cafe or restaurant.
+The offer is delicious pancakes for only $10.
+The website for more information is www.pancakehouse.com.
 |
+| posts | Create a happy and joyful image of a pregnant woman enjoying a relaxing day by the seacoast. The text emphasizes the excitement and love for the baby growing inside. |
 | covers& headers | Create a Facebook cover image to announce a flower exhibition event. |
-| marketing | Create a Thanksgiving gift certificate with an elegant and sophisticated design that offers a 50% discount on decorations. The message wishes a happy Thanksgiving to the recipient. |
-| marketing | Design a colorful store logo with an illustration of leaves to represent a fresh and vibrant shopping experience. The store may be named TB. |
-| marketing | Create a coupon that promotes a garden design service offered by Yard Culture.
-The coupon is a gift voucher valued at $49 and can only be used once.
-It is valid until November 15th.
-The coupon number is No:7496-4587.
+| posts | Design a Facebook post promoting a summer collection for women. Highlight the new arrival and offer a 60% discount for this week only. Encourage viewers to shop now. |
+| posts | Design a Facebook post advertising an organic cosmetics kit with natural and organic ingredients. The kit is a must-have for skincare and wellness. |
+| posts | Promote a cheesy Italian pizza with a discount offer of 30% for the first order. Encourage customers to visit the restaurant or order online. The promotion is for a limited time only. |
+| covers& headers | Design an advertisement for a crowdfunding platform called LocalBless.
+Encourage users to contribute to social projects by visiting the website and donating.
+Create a visually appealing image with a stone pattern background and triangles to convey a sense of stability and trust.
 |
-| marketing | Design a poster to promote a charity bike ride event to raise funds and support for a humanitarian cause.
-The event will be held on June 26th and the website for registration and donation is www.funbike.com.
+| posts | Design a Facebook post promoting a two-week summer camp that provides hands-on, project-based instruction in tech. The camp will take place from 12-14 August. |
+| posts | Create an Instagram story frame with a futuristic and fantasy-like design that features a man in an astronaut helmet and a space-themed background.
+The text #ThisIsMetaLife and SPACEDOUT are included.
 |
+| posts | Create a minimalistic and modern home design discount offer for a website or social media post.
+Encourage users to visit the website www.dreamhome.com to enjoy a 10% discount on their first design.
+|
+| covers& headers | Create an attractive YouTube thumbnail for a makeup review video featuring an African American woman.
+The video will showcase the woman’s favorite makeup products and provide a weekly review.
+|
+| posts | Promote a personal styling service for summer vacation outfits on Instagram. The contact number is +00 1357 2468 and website is www.website.com. |
 | marketing | Create an advertisement poster for a college merchandise store that has been operating since 1991.
@@
 |
-| posts | Create an Instagram story frame with a futuristic and fantasy-like design that features a man in an astronaut helmet and a space-themed background.
-The text #ThisIsMetaLife and SPACEDOUT are included.
+| posts | Create an Instagram story ad for an antique shop, promoting the sale of a timeless typewriter. Encourage customers to visit the store by swiping up on the story. |
+| covers& headers | Create a Youtube thumbnail for a video about real estate investments, providing tips from professionals. |
+| posts | Create a Facebook post promoting USA Independence Day tours with a patriotic theme.
+Encourage people to find their freedom in life by traveling more.
+Mention the offer of free cancellation.
+Include the website patrioutus.com for more information.
 |
-| posts | Create a recruiting Instagram post to advertise a job vacancy for a UI/UX designer. Encourage passionate individuals to apply and provide a link to apply in the bio. |
-| posts | Promote a personal styling service for summer vacation outfits on Instagram. The contact number is +00 1357 2468 and website is www.website.com. |
-| posts | Create an Instagram story promoting a special Christmas offer for a chocolate drink. The offer includes a 20% discount and encourages customers to order now. |
-| posts | Create an Instagram story ad for an antique shop, promoting the sale of a timeless typewriter. Encourage customers to visit the store by swiping up on the story. |
-| events | Design a holiday-themed gift card to promote the sale of gift boxes and packaging for the Christmas in July event. Include the website address www.yourbrand.com and store name. |
 | events | Create a Valentine’s Day greeting card with a cute and romantic design. The card includes the phrase ’Happy Valentine’s Day’ and ’Much love to you!’ |
-| events | Create a business card for a car service company that highlights their commitment to providing reliable and efficient car maintenance and repair services.
-The slogan ’Give your car wings’ emphasizes the quality of service provided.
-|
-| events | Design an advertisement card to promote the sale of premium quality school equipment, with a focus on enhancing the learning experience for students. |
-| events | Create a Women’s Day greeting card with signs of female gender, celebrating women and their achievements.
-The design should be beautiful and include floral elements.
-The card should be suitable for store use and have a message of ’Happy Women’s Day’.
-|
-| advertising | Create an Instagram ad for a real estate agency highlighting the modern and prestigious skyscraper building they have on offer.
-Encourage viewers to ’reach new highs’ by visiting the website at www.togalim.eu.
-Include contact phone numbers 126-3764-1376 and 301-17361-5085 for interested clients.
-|
-| advertising | Create an Instagram ad for a real estate agency highlighting the modern and prestigious skyscraper building they have on offer.
-Encourage viewers to ’reach new highs’ by visiting the website at www.togalim.eu.
-Include contact phone numbers 126-3764-1376 and 301-17361-5085 for interested clients.
-|
-| advertising | Create a Facebook advertisement for a store selling natural essential oils that promote healing for the body and soul.
-Highlight the use of floral and plant ingredients in the product.
-Encourage viewers to visit the store.
-|
-| advertising | Design a weekend sale advertisement for skincare products with a focus on promoting the cosmetic serum. All items are discounted up to 30% for this season. |
-| advertising | Create an Instagram ad for a camping festival and hiking event.
-The ad should promote the event and provide information about registration, including a website.
-Additionally, offer tips for surviving while hiking.
-|
-| advertising | Create an Instagram ad for a toy store’s Black Friday and Cyber Monday sale with a focus on the discount of 45% off on selected toys. |
-| creative | Design an attractive poster for a Mars botanical garden tour.
-Illustrate a greenhouse with a variety of extraterrestrial plants.
-Use vibrant colors to depict the potential biodiversity on Mars.
-|
-| creative | Create an enticing poster for a classical orchestra concert in the heart of the Amazon Rainforest.
-Depict the harmony between the music and the wilderness with a blend of vibrant greens and subtle sepia tones.
-Incorporate subtle hints of wildlife and flora within the musical instruments.
-|
-| creative | Design a dramatic poster for an opera performance in a natural Amazon Rainforest amphitheater.
-Use a bold, Baroque-inspired style to reflect the grandeur of opera, juxtaposed with the raw beauty of the forest.
-Show a dramatic scene unfolding on the forest stage.
-|
-| creative | Create a poster for an imaginary music festival where classical composers like Mozart and Beethoven are headliners.
-The design should merge classical elements with a modern music festival vibe, including stage setups and audience.
-|
-| creative | Construct an eye-catching poster for an upcoming academic conference on Quantum Physics.
-The design should encapsulate the mysterious and profound nature of the subject.
-Use a black hole or quantum particles as the main design element, and apply a deep space color scheme to create an intriguing atmosphere.
-Make sure the conference details and keynote speakers are clearly visible.
-|
-| Category |
-Image Generate Prompt |
-| marketing | Gift Certificate.
-blank, empty, layout, tb, certificate, offer, illustration, pattern, floral, flowers, pink, frame, sale.
-The image features a white background with a pink border.
-The border is adorned with a decorative floral pattern, giving the image a visually appealing and elegant appearance.
-The combination of the white background and the pink border with the floral pattern creates a harmonious and attractive design.
-|
-| posts | Instagram Story.
-astronaut, cosmic, cosmonaut, helmet, orange, science, space, spaceman, universe, story, instagram, ig, insta, frame.
-The image features a man wearing a spacesuit, sitting on a planet with his legs crossed.
-He is positioned in the center of the scene, and the planet is displayed in the background.
-The man appears to be a spaceman, possibly an astronaut, as he is dressed in a suit designed for space travel.
-The scene is set against a dark blue sky, which adds to the overall atmosphere of the image.
-|
-| marketing | Gift Certificate.
-agency, holiday, journey, service, tourism, travel, traveling, trip, vacation, tb, certificate, sale, discount, insurance, illustration, cartoon, blue.
-The image features a map with a variety of travel-related items scattered around it.
-There are several pairs of shoes, including a couple of boots, placed in different positions on the map.
-A hat is also present, positioned near the top left corner of the map.
-In addition to the shoes and hat, there are multiple books scattered across the map, with some located near the bottom left corner and others in the middle and top right areas.
-A pair of scissors can be seen in the top right corner of the map, and a clock is placed near the center of the map.
-|
-| events | Business Card European.
-data, card, health, home, professional, service, technology, icon, house, nursing, care, modern, medical, health care, contact, information.
-The image features a house with a pink roof, which is situated next to a street.
-The house is surrounded by a white fence, and there is a person standing on the sidewalk in front of the house.
-The person appears to be looking at the house, possibly admiring its architecture or considering it for purchase.
-The scene also includes a car parked on the street, adding to the overall atmosphere of the neighborhood.
-|
-| events | Instagram Post.
-wedding, dinner, rehearsal, banquet, event, holiday, celebration, marriage, food, party, restaurant, ceremony, leaf, jungle, tropical, exotic, brown, instagram, post, insta, ig, tb.
-The image features a brown background with a few green leaves scattered throughout the scene.
-The leaves are of various sizes and are positioned at different angles, creating a sense of depth and texture.
-The overall composition of the image is simple, with the brown background and green leaves being the main focus.
-|
-| events | Instagram Post.
-empowerment, equality, feminism, holiday, march, march 8, rights, women, women’s day, womensday, tb, illustration, female, girls, celebration, beautiful, greeting, sisterhood, instagram, insta, ig, post, flowers, floral.
-The image features two people standing next to each other, each holding a heart.
-The hearts are positioned above their heads, creating a visually striking scene.
-The people are standing in front of a backdrop of flowers, adding a touch of color and beauty to the image.
-|
-| covers& headers | Ticket DIN large.
-festival, art, abstract, geometric, graphic, illustration, pattern, shape, simple, visual, colorful, blue, ticket.
-The image features a colorful and abstract design, with a predominantly blue background.
-The design consists of various geometric shapes, including triangles, squares, and circles, arranged in a visually appealing manner.
-The shapes are positioned throughout the image, with some closer to the foreground and others further in the background.
-The overall composition creates a sense of depth and complexity, making the image visually engaging and interesting.
-|
-| advertising | Square Graphic Post.
-tb, instagram, insta, ig, post, style, fashion, trendy, fashionable, stylish, beautiful, clothes, model, woman, posing, modern, female, young, casual, beauty, attractive, look, color, modeling, lady, people, outfit, order, new arrival, pink, customer, testimonial, review.
-The image features a pink background with a white square placed in the center.
-The square is surrounded by a pink hue, giving it a slightly pinkish tint.
-The background is predominantly pink, with no other objects or distractions visible in the scene.
-|
+| posts | Create a Facebook post to celebrate International Pet Day with a cute little girl and a fluffy rabbit. |
 | Setting | Design LLM | Text-to-Background DM | Text-to-Object DM | Typography LMM | Reflect LMM | ||
@@
 This suitability arises from their ability to create backgrounds that allow for seamless incorporation of these various elements, fostering enhanced compositional flexibility.
+Typography LMM
++ Reflect LMM
+Typography LMM
++ Reflect LMM
 Typography LMM & Reflect LMM In Figure 19, we report the localization accuracy comparison results, such as mIoU, AP, AP, and AP, for the predicted text boxes based on the Typography LMM and the refined text boxes derived from the Reflect LMM.
@@
 This highlights our model’s superior performance, especially in terms of adherence to user intent and in the precision and readability of text rendering, as compared to DALLE3.
-E. Hyperparameters
+F Hyperparameters
 Table 10 outlines all the comprehensive training hyperparameters for our Cole system, which is composed of a hierarchical series of five generative models.
 The table is segmented into three parts: training settings, model settings, and data settings, listed in that order from top to bottom.
-Typography LMM
-+ Reflect LMM
-Typography LMM
-+ Reflect LMM
 Ours